### v1

# Synergy as the failure of distributivity

Abstract
A physical system is synergistic if it cannot be reduced to its constituents. Intuitively this is paraphrased into the common statement that ’the whole is greater than the sum of its parts’. In this manner, many basic parts in combination may give rise to some unexpected collective behavior. A paradigmatic example of such phenomenon is information. Several sources, which are already known individually, may provide some new knowledge when joined together. Here we take the trivial case of discrete random variables and explore whether and how it is possible get more information out of lesser parts. Our approach is inspired by set theory as the fundamental description of part-whole relations. If taken unaltered, synergistic behavior is forbidden by the set theoretical axioms. Indeed, the union of sets cannot contain extra elements not found in any particular set. However, random variables are not a perfect analogy of sets. We formalise the distinction, finding a single broken axiom - union/intersection distributivity. Nevertheless, it remains possible to describe information using Venn-type diagrams. We directly connect the existence of synergy to the failure of distributivity for random variables. When compared to the partial information decomposition framework (PID), our technique fully reproduces previous results while resolving the self-contradictions that plagued them and providing additional constraints on the solutions. This opens the way towards quantifying emergence in large systems.
Keywords emergence  $\cdot$ synergy  $\cdot$ information diagrams  $\cdot$ non-distributive  $\cdot$ parts and whole
1 Introduction
Synergism is loosely defined as “a whole that is greater than the sum of its parts”. Probability theory may be used to pour quantitative meaning into this equation-like definition. First, correlated random variables can capture the interaction between parts that is believed to lead to synergism. Second, information theory suggests entropy as a natural measure for random variables [1]. Alas, a basic property of the entropy of the joint distribution of random variables $X_{1}$ to $X_{N}$ is:
$H(X_{1},X_{2},\dots,X_{N})\leq\sum_{i=1}^{N}H(X_{i}),$ (1)
implying that the whole can never exceed the sum of its parts. Moreover, the left-hand side of this equation is maximized, not when the variables display interesting interactions like one would expect of a synergistic system but, rather, when they are all independent. A possible resolution comes from the concept of mutual information which, unlike entropy, is not subadditive.
$I((X,Y);Z)\lesseqqgtr I(X;Z)+I(Y;Z)$ (2)
Hence, if we do not consider the full information carried by $X$ and $Y$ but, rather, restrict ourselves to information that teaches us about a third variable, $Z$, then we can hope to find wholes that are larger than the sum of their parts. This peculiar property of mutual information became the starting point for study of information decompositions [2, 3, 4, 5] as the basis for the study of emergence. In their seminal paper, Williams and Beer [6] suggested that the Venn diagram often used to summarize the information measures associated with two correlated variables (Fig. 1) is more than a mere illustration. In fact, they suggested, the different regions in this diagram correspond to physical information atoms of the partial information decomposition (PID). For random source variables $X$, and $Y$ and relevant information $Z$ the Venn diagram corresponds to the following equations:
$I(X;Z)=R+U_{X},$ (3)
$I(Y;Z)=R+U_{Y},$ (4)
$I((X,Y);Z)=R+U_{X}+U_{Y}+S,$ (5)
where $R$ (the redundant information regarding $Z$ present in both source variables), $S$ (synergistic information obtained only when considering both source variables), and $U_{X}$, $U_{Y}$ (unique information in each of the source variables) are all non-negative. It turns out that this description, with the analogies it draws between set theory and random variables, has two significant drawbacks. First, the system of linear equations (3) is underdetermined and requires additional postulates to fix the extra degree of freedom. Indeed multiple redundancy measures [7, 8, 9, 10] cast doubt on the uniqueness of a Venn diagram approach. The second drawback is even more profound as it questions the very existence of a Venn diagram description. Namely, the axioms formulated to ensure the physical sense behind these definitions [6, 9] were shown to be self-contradictory when the number of random variables exceeds three [11]. This started a discussion regarding the validity of some information properties, such as non-negativity [12, 13] and moved the PID approach further from its set-theoretic origin.
In what follows, we illustrate how the self-contradiction mentioned above can find resolution when considering a fundamental distinction between set theory and probability theory: random variables, unlike sets, do not adhere to the distributivity axiom. This leads us to study a distributivity-free variant of set theory as a possible self-consistent theory of information atoms. Within this framework, we demonstrate that the presence of synergistic properties is a direct consequence of the broken axiom. This is fully consistent with the principle of conservation in the sense of (1), i.e. synergistic information does not appear from nowhere, but is still a part of the full information entropy. Furthermore, in the case of $N=3$ random variables we show that the amount of synergy quantifies the extent to which distributivity is breached. The acquired understanding allows us to fix long-standing discrepancies of PID [14, 11], introduces new constraints on redundancy measures, and provides tools for estimating and describing correlations between random variables.
2 Set-theoretic approach to information
Figure 1: A diagram representing different types of mutual information in the system of two random variables $X,Y$ wrt $Z$ by the part-whole relations. Redundant information is shared between $X$ and $Y$, unique information is a part of just one of them, while synergistic information is something that is only contained in the joint distribution, but not individual sources on their own.
In this section, we introduce the links between non-distributive sets, random variables, and synergy. We demonstrate these using the primary example of synergy: the logical XOR gate. A more general discussion will be presented in the next section.
2.1 Basic random variable operations
We start off by establishing some basic relations between information quantities and set theory. As mentioned before, the amount of information stored in a random variable $X$ is its Shannon entropy $H(X)$. It serves as a measure on the space of random variables. The set-theoretic union has an apparent equivalent - the operation of taking a joint distribution.
$X\cup Y=(X,Y)$ (6)
Consequently, the notion of a subvariable may be naturally defined as
$X\subset Y\Leftrightarrow\exists f:X=f(Y)\Leftrightarrow\exists Z:X\cup Z=Y,$ (7)
where $f$ is an arbitrary deterministic function.
Unfortunately, there is no well-defined counterpart to the intersection operator. It is common to view the mutual information function
$I(X;Y)=H(X)+H(Y)-H(X,Y)$ (8)
as the size of shared information between two variables. Indeed, (8) coincides with the inclusion-exclusion formula for two sets, which we would assume to hold in our setup. However, in the general case there is no mutual subvariable of $X$ and $Y$ whose entropy is equal to $I(X;Y)$ (Lemma 1). Nonetheless, a physically sensible ’intersection’ may be inferred in several trivial cases:
$H(X\cup Y)=H(X)+H(Y)\Leftrightarrow X\cap Y=\emptyset,$ (9)
$X\subset Y\Leftrightarrow X\cap Y=X$ (10)
2.2 Subdistributivity
Our set-theoretic intuition breaks down even further if we consider the XOR gate with two independent random inputs $O_{1},O_{2}$ and output $O_{3}$
$O_{1,2,3}=\begin{cases}0,&50\%\\ 1,&50\%\end{cases},$ (11)
$O_{3}=O_{1}\oplus O_{2}$ (12)
The pairwise independence between any two of its variables dictates that
$O_{1}\cap O_{2}=O_{2}\cap O_{3}=O_{1}\cap O_{3}=\emptyset$ (13)
At the same time, by definition the output is a deterministic function of the joint distribution of inputs, hence
$O_{3}\subset(O_{1}\cup O_{2})\Leftrightarrow O_{3}\cap(O_{1}\cup O_{2})=O_{3}$ (14)
Taken together, these violate the set-theoretic axiom of distributivity,
$O_{3}\cap(O_{1}\cup O_{2})=O_{3}\neq\emptyset=(O_{3}\cap O_{1})\cup(O_{3}\cap O_{2})$ (15)
Nevertheless, it can be shown that a weaker relation of subdistributivity holds for any three random variables (Lemma 2)
$(X\cap Z)\cup(Y\cap Z)\subset(X\cup Y)\cap Z$ (16)
Even though it is evident that random variables are quite different from sets, we argue that some of the logic behind PID may be recovered by extending set-theoretic notions, such as the inclusion-exclusion principle and Venn diagrams, to non-distributive systems.
2.3 Inclusion-exclusion formulas
There are other well-known examples of systems, whose operations are deemed analogous to set-theoretic, yet break the distributivity axiom. One example are vector spaces with the dimension measure and the direct sum and intersection operations. Note that, if we consider only two vector spaces, the inclusion-exclusion formula holds. Similarly, the inclusion-exclusion principle holds for the case of only two random variables. This is since its proof does not rely on the distributivity axiom. When the intersection between $X$ and $Y$ exists, its entropy satisfies
$H(X\cup Y)=H(X)+H(Y)-H(X\cap Y)$ (17)
Beyond two variables, we encounter some new properties. For example, the inclusion-exclusion formulas for the XOR gate can be obtained by repeatedly applying the 2-variable formula (17) while setting all pairwise intersections to zero
$H(O_{1}\cup O_{2}\cup O_{3})=H(O_{1}\cup O_{2})+H(O_{3})-H((O_{1}\cup O_{2})\cap O_{3})=$ (18)
$=H(O_{1})+H(O_{2})+H(O_{3})-H((O_{1}\cup O_{2})\cap O_{3})$ (19)
This formula disagrees with the analogous set-theoretic formula (for non-intersecting sets) only in the last term, which is non-zero precisely due to the subdistributivity. Note that while the rest of the terms are symmetric with respect to permutation of indices, the expression $(O_{1}\cup O_{2})\cap O_{3}$ is not as it explicitly depends on the order of derivation. This essentially leads to three different inclusion-exclusion formulas. Nevertheless, the size of this term remains invariant
$H((O_{1}\cup O_{2})\cap O_{3})=H((O_{1}\cup O_{3})\cap O_{2})=H((O_{2}\cup O_{3})\cap O_{1})$ (20)
In our case, the distributivity-breaking expressions may be simplified using (14).
$(O_{i}\cup O_{j})\cap O_{k}=O_{k}$ (21)
2.4 Construction of Venn-type diagram for the XOR gate
The non-uniqueness of inclusion-exclusion formulas complicates the construction of Venn-type diagrams in subdistributive systems. A way of tackling this and well as some intuition can be traced via our XOR gate example.
In set theory, Venn diagrams act as graphical representations of the inclusion-exclusion principle. The inclusion-exclusion formula sums up the size of all possible intersections between the participating sets. For correct bookkeeping, this is done with alternating signs that account for the covering number - the number of times each intersection is ’counted’ as a part of some set. In classical set theory, the covering number of an intersection is trivially the number of sets which are being intersected. However, (18), includes the distributivity-breaking term which is absent from this classical theory and whose covering number is not evident. This term appears with a negative sign which signifies an even-times covered region. In this three variable system, the only even, non-empty alternative is a $2$-covered region. From another perspective, in each of the three possible formulas $O_{k}$ is covered once by itself and one more time by the union $O_{i}\cup O_{j}$ (though not by $O_{i}$ or $O_{j}$ individually). As for the size of this region, independent of $k$ and therefore of the order of decomposition, it measures at $1$ bit of information. Denoting this area as $\Pi_{s}$, we have
$\Pi_{s}[2]\equiv H((O_{i}\cup O_{j})\cap O_{k})=1\text{ bit},$ (22)
where the covering number is indicated in the brackets [][][ ]. The region $\Pi_{s}$ does not contain enough information to describe the whole system, as the total amount of entropy in the XOR gate is
$H(O_{1}\cup O_{2}\cup O_{3})=2\text{ bit}$ (23)
To draw the diagram, we need to find the covering of the remaining $2-1=1$ bit region/regions. This may be accomplished by borrowing two properties of set-theoretic diagrams.
First of all, in a system of $N$ arbitrary random variables $X_{1},\dots X_{N}$ the total entropy of the system must be equal to the sum of all diagram regions $\Pi_{i}[c_{i}]$
$H(X_{1},\dots X_{N})=\sum_{i}\Pi_{i}[c_{i}]$ (24)
Second, the sum of individual variables’ entropies is equal to the sum of region sizes times their corresponding covering numbers $c_{i}$
$\sum_{k=1}^{N}H(X_{k})=\sum_{i}c_{i}\Pi_{i}[c_{i}]$ (25)
These properties provide straightforward physical meaning to the information conservation law: adding new sources should either introduce new information or increase the covering of existing regions. Information cannot spontaneously arise nor vanish.
Let us assume that in addition to $\Pi_{s}$ the diagram of the XOR gate contains several more regions $\Pi_{i}$. To calculate their sizes and coverings we apply (24-25)
$\sum_{i}(c_{i}-1)\Pi_{i}[c_{i}]=0\text{ bit}$ (26)
We use the fact that information is non-negative and discard meaningless empty regions. The above equation then allows for a single $1$-bit region, which is covered once
$\Pi_{g}[1]=1\text{ bit}$ (27)
Similarly to (3), we have now decomposed the information inside the XOR gate variables into several basic parts. As we will soon see, these are nothing else but the information atoms (a more general set of atoms, which will include all previously found by Williams and Beer). To respect the physical meaning behind them, we demand the structure of the resulting Venn diagram to be well-defined. In other words, despite the existence of three different versions of inclusion-exclusion formula (18), they are all assumed to describe the same universal decomposition. Indeed, our result remains invariant with respect to index permutations in terms of region (atom) sizes and covering numbers.
We will now use this universality assumption to determine the part-whole relations, i.e. the shape of the Venn diagram. (21,22) dictate that the atom $\Pi_{s}$ corresponds to all variables $O_{k=\overline{1,3}}$ at the same time
$\Pi_{s}=H(O_{1})=H(O_{2})=H(O_{3})$ (28)
It can therefore be thought of as a $2$-covered triple intersection between $O_{1},O_{2}$ and $O_{3}$. This a drastic divergence from classical classical set theory where an intersection between $n$ sets is covered only exactly $n$ times. As we shall see, without distributivity, $n$ variables can have multiple intersection regions with different covering numbers $c\leq n$.
Moving on to the second atom in this system: $\Pi_{g}$ only appears as a leftover from taking the difference between the whole system and $\Pi_{s}$ and by set-theoretic intuition it does not intersect with $O_{k}$ for any $k$. As such it is not a part of any single variable. We can now combine all our findings into a single system of decomposition equations
$H(O_{1})=\Pi_{s},$ (29)
$H(O_{2})=\Pi_{s},$ (30)
$H(O_{3})=\Pi_{s},$ (31)
$H(O_{1}\cup O_{2}\cup O_{3})=\Pi_{s}+\Pi_{g},$ (32)
where one can see that any double joint distribution $O_{i}\cup O_{j\neq i}$ is equal to the whole system $O_{1}\cup O_{2}\cup O_{3}$. This system generates the Venn-type diagram that represents the information distribution inside the XOR gate - Fig. 2.
Figure 2: A Venn-type diagram for the XOR gate. Each variable is represented by a primary color circle (red, yellow, blue) while the outer circle outlines the whole system. Of the total $2$ bits of the XOR gate, one is covered two times and is represented by the inner disk. Since it is covered twice this area is colored by pairwise color-blends (orange, purple, and green). Since it is covered by three variables it includes patches of all three possible blends. A critical difference between this diagram and a set-theoretic one is that even though the three variables have no pairwise intersections, the inner disk representing the ’mutual’ content of all three variables is non-empty. The remaining $1$ bit is covered once and resides only inside the joint distribution. Since this area is covered once, it is colored by primary colors. Patches of all three colors are used since this area in a part of any single variable.
2.5 Synergy as an information atom
Let us now compare the results of set-theoretic approach to the original PID construction. The XOR gate is a paradigmatic example of synergistic behaviour. Due to its degenerate nature (lack of intersections between variables) its PID is unique, can be calculated explicitly, and is known to contain only a single bit of synergistic information. Indeed, treating the variables $O_{1},O_{2}$ as inputs and $O_{3}$ as the output into the PID equations (3), we get
$R=U_{1}=U_{2}=0\text{ bit},$ (33)
$S=1\text{ bit}$ (34)
On the other hand, the left side of each line in (3) may be rewritten using our set-theoretic formalism by definition as intersections of random variables
$I(X;Z)=H(X\cap Z),$ (35)
$I(Y;Z)=H(Y\cap Z),$ (36)
$I((X,Y);Z)=H((X\cup Y)\cap Z)$ (37)
For the XOR gate all of these are well-defined: the former two are empty, while the last line corresponds to a specific part of our diagram. It produces a single non-trivial relation between the original definition of synergistic information and the quantities we have derived
$I((O_{1},O_{2});O_{3})=H((O_{1}\cup O_{2})\cap O_{3})=S$ (38)
In fact, the non-set-theoretic term of the inclusion-exclusion formula (18) and, consequently, the atom $\Pi_{s}$ are equal to the synergistic information
$\Pi_{s}=S$ (39)
This is a trivial case of our main result: in a trivariate system synergy measures the breaking of distributivity in the system. Besides, it does not contradict the subadditivity of entropy in any way. The synergistic information is not new to the system and is a part of the sources’ full information content. It is actually redundant in the sense that it appears two times ($\Pi_{s}[2]$ is $2$-covered).
The nature of ghost atom $\Pi_{g}$ in the diagram is deeply connected to this outcome even though it does not explicitly participate in the PID. Consider the individual contributions by each of the sources $i=1,2$ measured via the mutual information function
$I(O_{i};O_{3})=H(O_{i})+H(O_{3})-H(O_{i}\cup O_{3})$ (40)
Using (29) we can rewrite this in terms of atoms
$I(O_{i};O_{3})=\Pi_{s}+\Pi_{s}-(\Pi_{s}+\Pi_{g})=\Pi_{s}-\Pi_{g}=0$ (41)
The equality between the synergistic and ghost atoms ensures that the former is exactly cancelled from the individual contribution by each source. Synergistic information is, of course, still present in the ’whole’ (38). This circumstance is responsible for creating the illusion of synergy appearing out of nowhere.
3 General trivariate decomposition
In this section, we explore the properties of the non-synergistic parts of the diagram. In combination with the preceding discussion about the nature of synergy in the XOR gate, this will allow us to formulate a complete description of the arbitrary trivariate system.
3.1 Reintroducing redundant information
The XOR gate analyzed above provides a simple example for synergism but includes no redundant information. As a first step towards an general trivariate variable we append more variables to the XOR setup. This will produce diagrams that contain all PID atoms. Let the random variables $R,U_{1},U_{2}$ be mutually independent among themselves as well as with respect to the XOR variables $O_{1,2,3}$. Now consider the system of three variables
$X_{1}=R\cup U_{1},$ (42)
$X_{2}=R\cup U_{2},$ (43)
$X_{3}=R\cup U_{1}\cup U_{2}$ (44)
The new parts correspond to redundant information of different kinds, as evidenced by their presence in pairwise intersections
$X_{1}\cap X_{2}\cap X_{3}=R,$ (45)
$X_{1}\cap X_{3}=R\cup U_{1},$ (46)
$X_{2}\cap X_{3}=R\cup U_{2}$ (47)
When considered on their own, they form an ordinary Venn diagram with triple intersection $R$ and two double intersections $U_{1},U_{2}$
$H(X_{1})=H(R)+H(U_{1}),$ (48)
$H(X_{2})=H(R)+H(U_{2}),$ (49)
$H(X_{3})=H(R)+H(U_{1})+H(U_{2}),$ (50)
$H(X_{1}\cup X_{2}\cup X_{3})=H(X_{i}\cup X_{j\neq i})=H(R)+H(U_{1})+H(U_{2})$ (51)
Changing the variables to $Y_{i}=X_{i}\cup O_{i}$, we get a superposition of systems (29, 48). The end result as summarized in (Fig. 3) is simply a combination of the corresponding diagrams. In the context of PID equations it coincides with Williams and Beer’s construction for inputs $Y_{1,2}$ and output $Y_{3}$
$I(Y_{1};Y_{3})=H(R)+H(U_{1}),$ (52)
$I(Y_{2};Y_{3})=H(R)+H(U_{2}),$ (53)
$I((Y_{1},Y_{2});Y_{3})=H(R)+H(U_{1})+H(U_{2})+\Pi_{s}$ (54)
This result is not too practical per se, since the intersections (45) are required to coincide with some independent random variables and the chosen synergy is obviously constrained to $1$ bit. However, in the following we prove that the diagram of this form describes the general structure of an arbitrary $3$-variable decomposition.
Figure 3: Sample diagram containing both synergistic and redundant parts. $Y_{1}$ and $Y_{2}$ are represented by red and yellow circles while $Y_{3}$ (blue boundary) is the union of these two circles (see section 2A). The set-theoretic atoms $R,U_{1,2}$ are correspondingly $3$- and $2$-covered. Being a part of the normal set-theoretic diagram these atoms are colored by the plain three and two color blends. Note that while the synergistic atom $\Pi_{s}$ is drawn as a part of the central region, it is not a part of any double intersection $Y_{3}\cap Y_{1,2}$ (see Fig. 2).
3.2 Extended random variable space
Lack of a proper description for ’information intersections’ severely limits our ability to decompose the information content of more general random variable systems. Our solution for this issue is inspired by an elegant duality between set theory and information quantities found by Hu in [15] and further elaborated in [16]. It simply extends the space of random variables to include all elements produced by the operations $\cup$,$\cap$ , ∩ (6-9). The entropy is extended as a (non-negative) measure $\hat{H}$ with two simple conditions
$\hat{H}(X)=0\Leftrightarrow X=\emptyset,$ (55)
$X\cap Y=\emptyset\Leftrightarrow\hat{H}(X\cup Y)=\hat{H}(X)+\hat{H}(Y)$ (56)
[…]
$\Delta\cap Y=\emptyset,$ (57)
$\Delta\cup(X\cap Y)=X$ (58)
Note that it is not a function, i.e the result may not be unique (114-116).
A complete description of information intersections will allow us to not only find the decomposition for a system of $3$ arbitrary random variables, but also extend our theory to a multivariate setup. To approach the problem of characterizing information atoms in the $3$-variable case, we need a general trivariate inclusion-exclusion formula. As stated previously, the bivariate version (17) holds without alterations (Lemma 3). Now, in contrast, we get a familiar distributivity-breaking term, which depends on the derivation (Theorem 1)
$\hat{H}(X_{1}\cup X_{2}\cup X_{3})=\hat{H}(X_{1})+\hat{H}(X_{2})+\hat{H}(X_{3})-$ (59)
$-\hat{H}(X_{1}\cap X_{2})-\hat{H}(X_{1}\cap X_{3})-\hat{H}(X_{2}\cap X_{3})+\hat{H}(X_{1}\cap X_{2}\cap X_{3})-\Delta\hat{H},$ (60)
$\Delta\hat{H}=\hat{H}(((X_{i}\cup X_{j})\cap X_{k})\backslash((X_{i}\cap X_{k})\cup(X_{j}\cap X_{k})))$ (61)
$\forall i\neq j\neq k$ (62)
A visualisation for one possible variant of this formula can be seen on Fig. 4.
Figure 4: A single realization of inclusion-exclusion principle for three variables. The new region, corresponding to the distributivity-breaking term is represented via a checkered pattern. Covering numbers are written for each sector and highlighted by the colors. This is not full a Venn-type diagram that defines the information atoms, hence its structure is clearly not invariant wrt variable permutations.
3.3 Set-theoretic solution
To study redundant information, we shall first consider a system without synergy. In this case distributivity axiom holds and the decomposition is equivalent to the Venn diagram for $3$ sets. To describe the atoms it is convenient to use the antichain notation known from the PID (146-150). Their sizes may be obtained by the means of Möbius inversion
$\forall i\neq j\neq k\text{ }\exists\Pi_{\{1\}\{2\}\{3\}}[3],\Pi_{\{i\}\{j\}}[2],\Pi_{\{i\}}[1]:$ (63)
$\Pi_{\{1\}\{2\}\{3\}}=\hat{H}(X_{1}\cap X_{2}\cap X_{3}),$ (64)
$\Pi_{\{i\}\{j\}}+\Pi_{\{1\}\{2\}\{3\}}=\hat{H}(X_{i}\cap X_{j}),$ (65)
$\Pi_{\{i\}}+\Pi_{\{i\}\{j\}}+\Pi_{\{i\}\{k\}}+\Pi_{\{1\}\{2\}\{3\}}=\hat{H}(X_{i})$ (66)
As we already know, not all systems may be described as distributive. To derive the necessary condition for the set-theoretic solution to be applicable, consider the $3^{\text{rd}}$ order interaction information function
$I_{3}(X_{1};X_{2};X_{3})=\sum_{i}H(X_{i})-\sum_{i\neq j}H(X_{i},X_{j})+H(X_{1},X_{2},X_{3})$ (67)
Importantly, this is exactly the difference between the ’sum of the parts’ and the ’whole’ as depicted in the two sides of (2). Using the full set of inclusion-exclusion formulas (17, 59) one can find that for any system it is equal to
$I_{3}(X_{1};X_{2};X_{3})=\hat{H}(X_{1}\cap X_{2}\cap X_{3})-\Delta\hat{H}$ (68)
Plugging in $\Delta\hat{H}=0$ by the requirement of distributivity, we find that for a set-theoretic system the interaction information is non-negative, i.e. the whole is always less than the sum of the parts. It immediately follows that under the same condition all atoms of decomposition (63) are non-negative (Lemma 9).
3.4 Arbitrary trivariate system
At this point we have studied two opposite cases of a completely synergistic system (XOR gate) and one without any synergy (set-theoretic solution). To describe three arbitrary variables, any general decomposition must be able to replicate both of them. To get some intuition behind the construction of such decomposition, consider the constraint on the interaction information derived in the previous section. Conversely, for the XOR gate $I_{3}$ is strictly negative
$I_{3}(O_{1};O_{2};O_{3})=-\Pi_{s}=-1\text{ bit}$ (69)
We will see that the decomposition, which combines set-theoretic and synergistic/ghost atoms, allows for any values of interaction information. Compared to section 2A, we are not tied to specific random variables in our description of information atoms. Let us simply add synergistic and ghost atoms of arbitrary (yet equal) sizes to the set-theoretic solution (Fig. 5)
$H(X_{i})=\Pi_{s}+\sum_{\text{set-theor. atoms}}\Pi,$ (70)
$H(X_{i},X_{j\neq i})=\Pi_{s}+\Pi_{g}+\sum_{\text{s.t. atoms}}\Pi,$ (71)
$H(X_{1},X_{2},X_{3})=\Pi_{s}+\Pi_{g}+\sum_{\text{s.t. atoms}}\Pi,$ (72)
$\Pi_{s}=\Pi_{g}$ (73)
The resulting system of equations (as presented in detail in 289) indeed provides a non-negative decomposition for the arbitrary trivariate system (Lemma 12). It becomes the minimal solution to the problem, containing the smallest set of necessary atoms. The interaction information is now a difference of two terms:
$I_{3}(X_{1};X_{2};X_{3})=\Pi_{\{1\}\{2\}\{3\}}-\Pi_{s}\lesseqqgtr 0$ (74)
Figure 5: A graphical illustration for the general solution of the trivariate problem. Compared to the Venn diagram for $3$ sets, two new regions here are the $2$-covered part of triple intersection $\Pi_{s}$ (synergistic atom) and a ghost atom $\Pi_{g}$, which is not a part of any single initial variable. Similarly to Fig. 4, colors indicate the coverings: 3 primary colors (red, yellow, blue or their checkered combination) correspond to 1-covered atoms, the overlay of any 2 colors (orange, purple, green or their checkered combination) is 2-covered and the overlay of all 3 colors (brown) is 3-covered.
To find the physical meaning behind the recovered solution, we once again choose $X_{3}$ as an output and reduce our attention to the mutual information between it and the individual inputs $X_{1},X_{2}$ or their joint distribution $X_{1}\cup X_{2}$. Only $4$ of the atoms from Fig. 5 appear in the corresponding PID equations
$I(X_{1};X_{3})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{\{1\}\{3\}},$ (75)
$I(X_{2};X_{3})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{\{2\}\{3\}},$ (76)
$I((X_{1},X_{2});X_{3})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{s}+\Pi_{\{1\}\{3\}}+\Pi_{\{2\}\{3\}}$ (77)
The result fully captures the structure behind Williams and Beer’s definitions
$\Pi_{\{1\}\{2\}\{3\}}\equiv\text{Redundancy},$ (78)
$\Pi_{\{1\}\{3\}}\equiv\text{Unique information in }X_{1},$ (79)
$\Pi_{\{2\}\{3\}}\equiv\text{Unique information in }X_{2},$ (80)
$\Pi_{s}\equiv\text{Synergy}$ (81)
The role of ghost atom $\Pi_{g}$ remains the same as discussed in section 1E. As mentioned in [6], any increase of synergy with respect to redundancy leads by (74) to the whole being greater than the sum of the parts in the sense of (2).
To summarize, we have shown how synergy naturally follows from set-theoretic arguments and does not contradict the conservation of information. Furthermore, comparing (68) and (74), we see that the amount of synergistic information is equal to the extra contribution of subdistributivity in the inclusion-exclusion principle
$S=\Pi_{s}=\Delta\hat{H}$ (82)
This allows us to conclude that synergistic behaviour is a direct consequence of distributivity-breaking. The above formula may be understood as the fundamental definition of synergy. Originating from the very axioms of set theory, it can be applied far beyond information.
Notice that system (289) includes $8$ equations for $9$ unknowns. Therefore, we can not uniquely determine the sizes of atoms. This problem is of the same nature as the ambiguity of PID atoms mentioned in the introduction (3). Using some of the extended random variable space properties we gain several insights on this issue. First, both redundant and synergistic information are symmetric with respect to variable permutations. They will be the same for all three possible PIDs with each variable taken as an output. Second, our formalism allows to formulate a new lower bound on the redundant information using a relation $\sim$ between different inputs (Lemma 15, Theorem 5)
$\max_{\begin{subarray}{c}X_{1}^{\prime}\sim X_{1}\\ X_{2}^{\prime}\sim X_{2}\end{subarray}}I_{3}(X_{1}^{\prime};X_{2}^{\prime};X_{3})\leq R,$ (83)
$X\sim X^{\prime}\Leftrightarrow\begin{cases}Y\rightarrow X\rightarrow X^{\prime}\\ Y\rightarrow X^{\prime}\rightarrow X\end{cases}$ (84)
4 Towards a multivariate information decomposition
In this section we lay the foundation for a consistent theory of the multivariate decomposition and resolve the contradictions between PID axioms.
4.1 Partitioning a non-distributive space
The first step would be to generalize the inclusion-exclusion principle for spaces without distributivity to $N$-variables. While for $N=3$ it remains invariant to permutations at least regarding the structure of the formula and sizes of its terms, for larger systems the lack of distributivity leads to a set of entirely different formulas. To obtain them we follow the same procedure as in set theory: gradually splitting unions into parts using the $2$-variable formula. We arrive at
$\hat{H}(X_{1}\cup X_{2}\cup\dots\cup X_{N})=$ (85)
$=\sum_{i}\hat{H}(X_{i})+\sum_{n=2}^{N}\sum_{i_{1},\dots,i_{n}}(-1)^{n-1}\hat{H}\left(\bigcap_{k=\overline{1,n}}X_{i_{k}}\right)+\sum(-1)^{C-1}\hat{H}(\Delta X[C]),$ (86)
where the first part matches the set-theoretic version, whereas the additional sum of distributivity-breaking differences $\Delta X$ depends on the derivation order and each term is characterized by a covering number $C$, which determines its sign. To see, how such definition naturally arises from set-theoretic intuition, consider a term appearing on an arbitrary step of the derivation
$\Xi[C]=\bigcap\left(\bigcup X_{i}\right)$ (87)
The covering number $C$ for it is defined trivially as the number of intersecting union-brackets in (87). To apply the $2$-variable inclusion-exclusion formula, we choose a single union-bracket and divide it into $2$ parts
$\Xi[C]=\left((X_{1}\cup\dots\cup X_{i})\cup(X_{i+1}\cup\dots\cup X_{m})\right)\cap\Xi^{\prime}$ (88)
For simplicity here indices are renamed to $\overline{1,m}$ and the rest of the expression is denoted by $\Xi^{\prime}$. We proceed in the same manner as during the proof of (Theorem 1). By subdistributivity axiom, we extract a difference term and break down the union into several parts
$\hat{H}(\Xi[C])=\hat{H}(\Delta X[C])+$ (89)
$+\hat{H}\left((X_{1}\cup\dots\cup X_{i})\cap\Xi^{\prime}\right)+$ (90)
$+\hat{H}\left((X_{i+1}\cup\dots\cup X_{m})\cap\Xi^{\prime}\right)-$ (91)
$-\hat{H}\left((X_{1}\cup\dots\cup X_{i})\cap(X_{i+1}\cup\dots\cup X_{m})\cap\Xi^{\prime}\right)$ (92)
Most of the terms have the same form as $\Xi$ and the procedure repeats for them. Similarly to the $3$-variable version, the covering of the difference term $\Delta X$ is assumed to be equal to $C$, which by design of the partitioning algorithm (87-89) coincides with its sign in (85). The process starts from the union of all variables and continues until one of the versions of (85) is reached, i.e. no union operations are left. Note that partitioning can be done in many ways leading to very different terms $\Xi$.
It will be important for our further discussion, that the terms $\Xi$ may be compared by a trivial order of inclusion, provided they appear in the same version of partitioning algorithm. The resulting lattices for each partition may then be merged into a single inclusion lattice $(L^{I},\preceq)$ by identifying the same nodes found in different partitions. Thus, $\Xi_{i}\preceq\Xi_{j}$ iff in at least one partition $\Xi_{i}\subset\Xi_{j}$. This lattice is in fact isomorphic to the antichain lattice (Theorem 2).
4.2 Information atoms based on part-whole relations
In order to keep the set-theoretic intuition, we once again assume that despite being different, the formulas of partitioning (85) all describe the same system and correspond to a single universal Venn-type diagram. This paradox may be resolved by elaborating on the notion of information atoms.
The information atoms - regions of the diagram - are defined as the basic pieces of information. Hence, all more complex quantities must be made up of atoms. The most general set of ’information pieces’ has been defined by us in the form of (87). We may therefore reformulate the problem of finding the decomposition as the problem of defining part-whole relations between the atoms and partition terms $\Xi$. A perfect tool for such purpose was proposed in [17] in the form of the parthood table. It is a matrix with entries 00 or $1$, which defines whether a given atom is a part of a particular $\Xi$
$\hat{H}(\Xi_{i}[C_{i}])=\sum_{j}f_{ij}\Pi_{j}[c_{j}],$ (93)
$f_{ij}=0,1$ (94)
The parthood table determines which atoms can exist in the system at all as part-whole configurations. It only depends on the number of variables
$f=f(N)$ (95)
The summands $\Pi$ are non-negative functions of the initial random variables and represent the actual sizes of atoms
$\Pi_{j}=\Pi_{j}(X_{1};X_{2};\dots X_{N})$ (96)
Similarly to the original scheme [17], we postulate the monotonicity, i.e. consistency with the inclusion lattice
$\Xi_{i}\preceq\Xi_{j}\Rightarrow\forall k\hookrightarrow f_{ik}\leq f_{jk}$ (97)
It relates the table’s entries within themselves by a simple rule: if one partition term is included in the other, all the atoms from the decomposition of former should be present in the decomposition of the latter. The assumption of universality explains the merging of inclusion lattices from different partitions into a single lattice $(L^{I},\preceq)$. Both partition terms $\Xi$ and atoms $\Pi$ describe the regions of a single Venn-type diagram, hence the relations between them must hold for every partition.
The covering numbers $c_{j}$ of the atoms can be derived from the coverings of partition terms $C_{i}$. In a set-theoretic Venn diagram an atom $\Pi_{j}$ cannot be a part of $\Xi_{i}$ with a higher covering $C_{i}>c_{j}$. Besides, each atom is a part of at least one partition term with the same covering $C=c$. The covering number is then fixed as a function of the parthood table
$c_{j}=\max_{i:f_{ij}=1}C_{i}$ (98)
The information conservation law (25) is the final condition that preserves the physical meaning of the covering numbers - the number of times the same information appears in the system.
4.3 Examples of completely solvable systems
The existence of a general solution for an N-variable system is not guaranteed. In the systems with large $N$ the number of unknowns dramatically increases compared to the number of equations on the entropies (85), which serve as boundary conditions for our measure $\hat{H}$. This means that if a solution does exist, it might not be unique. For a specific set of degenerate systems it is, however, still possible to calculate the sizes of all atoms. Then, using the additivity in the solution space (Lemma 8), we may determine the structure of the general decomposition by taking the superposition of known solutions similarly to how we got Fig. 5.
Set-theoretic solution for $N$ variables
If all distributivity-breaking differences $\Delta X$ are put to zero, there exists only a single possible partition, coinciding with the inclusion-exclusion principle for $N$ sets. The solution for this setup consists of $12^{N}-$ 1 atoms obtained via Möbius inversion (216-219). It is valid iff all atom sizes are non-negative, providing us with a criterion
$\forall i_{\overline{1,n}}=\overline{1,N}\hookrightarrow\sum_{m=n}^{N}(-1)^{m-n}\sum_{i_{n+1},\dots i_{m}}I_{m}(X_{i_{1}};\dots X_{i_{m}})\geq 0$ (99)
Set-theoretic systems do not exhibit synergistic properties. It can be shown that for these systems mutual information as a function of random variables also becomes subadditive (Lemma 10). The breaking of distributivity is therefore a necessary condition for the presence of synergy in multivariate systems.
XOR gate
We show that the solution found for the XOR gate is unique in the parthood table formalism (233-235, Theorem 3). This reinforces our proposal of synergistic and ghost atoms as physical entities and proves that Fig. 5 is the minimal solution for the trivariate problem. The resulting table (Table 1) also shows that the synergistic atom $\Pi[2]$ is indeed a part of each variable $X_{i},i=\overline{1,3}$. Moreover, using this example, we discover that it is impossible to associate atoms with any members of the extended random variable space (Lemma 11), leaving the question about their precise mathematical nature unanswered.
$N$-parity and other variations
Generalizing XOR gate to an arbitrary number of variables yields the $N$-parity setup, for which the solution has a similar form: a single $2$-covered synergistic atom $\Pi_{s}$ and a set of $N-2$ $1$-covered ghost atoms $\Pi_{g_{n}}$ of the same size (260-270)
$\Pi_{s}[2]=1\text{ bit},\qquad\Pi_{g_{n=\overline{1,N-2}}}[1]=1\text{ bit}$ (100)
It is also possible to get a more complex solution by including $m_{i}$ copies of each variable $X_{i}$ in the previous setup (271-288). The solution consists of the same atoms as the normal $N$-parity, yet the coverings now depend on the sizes of equivalence classes $m_{i}$. Without loss of generality ordering them by size we get
$m_{1}\geq m_{2}\geq\dots\geq m_{N},$ (101)
$\Pi_{s}[m_{1}+m_{N}]=1\text{ bit},\qquad\Pi_{g_{n=\overline{1,N-2}}}[m_{n+1}]=1\text{ bit}$ (102)
4.4 Resolving the PID self-contradiction
The existence of any multivariate PID was previously believed to be disproved [11] by employing a simple example that could not be solved without discarding one of the basic PID axioms (351-361). This example consists of 4 specific variables: the fair coins $O_{1},O_{2},O_{3}=O_{1}\oplus O_{2}$ from the XOR example act as inputs, while the output is their joint distribution $X_{4}=(O_{1},O_{2},O_{3})$. It was claimed that this system contains $3$ synergistic atoms: $\{1\}\{23\}$, $\{2\}\{13\}$ and $\{3\}\{12\}$. These were summed up to give three bits of information which is more than the total of two bits present in the entire system. They, therefore, conclude that the non-negativity of information is not respected.
One can easily check that our multivariate decomposition automatically satisfies the general PID axioms (Lemma 14). The problem of the past attempts, however, was not a consequence of the axioms themselves, but followed from an inaccuracy in the PID structure. Here we will locate the error and show that our approach is free from it. As before, the PID atoms are a subset of information atoms in our Venn-type diagram. In the $N$-variable PID they may be easily identified by considering the mutual information between the union of all inputs $X_{1},\dots X_{N}$ and the output $X_{N+1}$
$I(X_{1},\dots,X_{N};X_{N+1})=\hat{H}((X_{1}\cup\dots\cup X_{N})\cap X_{N+1})$ (103)
Accordingly, the atoms that are of our interest lie in the above intersection and are defined by the submatrix of the full parthood table
$f_{ij}:\Xi_{i}\preceq(X_{1}\cup X_{2}\cup\dots X_{N})\cap X_{N+1}$ (104)
In particular it may be shown that when the output is equal to the joint distribution of all inputs, then the entropy of inputs coincides with mutual information and hence all non-zero atoms from the full diagram appear in the PID (Lemma 13). The diagram itself can be easily obtained from the system of inputs $X_{1},\dots X_{N}$ alone by keeping atoms’ form and sizes ans simply adding one extra cover to each of them (Theorem 4 4). This is exactly the type of system that was used in [11]. Its decomposition mirrors the full diagram of the XOR gate with coverings increased by $1$.
$\Pi_{s}[3]=1\text{ bit},$ (105)
$\Pi_{g}[2]=1\text{ bit},$ (106)
$I(O_{i=\overline{1,3}};O_{4})=\Pi_{s},$ (107)
$I((O_{i},O_{j\neq i});O_{4})=I((O_{1},O_{2},O_{3});O_{4})=\Pi_{s}+\Pi_{g}$ (108)
We can see that in place of three synergistic atoms, there is only one symmetric atom $\Pi_{s}[3]$. The confusion that arose in [11] occurred since the different forms of the inclusion-exclusion principle were considered separately and it was assumed that each version would create its own synergistic atom.
5 Conclusion
Previous attempts for studying synergistic information using set-theoretic intuition have led to self-contradictions. In this work we point out that the non-distributivity of random variables corresponds to a well-defined variant of set-theory. This lifts previous contradictions and reopens this direction of research. We employ our results to construct a universal Venn-like diagram for an arbitrary $3$-variable system and demonstrate how synergism to be a direct consequence of distributivity breaking.
Our results do not fully solve the problem at hand. First, precise calculation of atom sizes was left unanswered and might require a more explicit description of information intersections. Another caveat is that although we constructed the equations that describe a self-consistent multivariate information decomposition, the existence of solution for $N$ arbitrary random variables is yet to be proven.
Nevertheless, this work lays the basis for a possible multivariate theory. The analysis we present reestablishes the concept of information decompositions as a foundation for further enquiry in quantifying emergence. In this context, information theory serves as a mere illustration: the mechanism we describe explains the nature of synergy using solely set-theoretic concepts and can be applied to any emergent physical system.
Appendix A Properties of the random variable (RV) space
Despite the existence of natural definitions for unions (joint distributions) and subsets (deterministic functions), the RV space does not contain elements that could be seen as analogues to set-theoretic intersections in all cases. The inclusion-exclusion principle is required to hold at least in the 2-variable case [lemma 3] and, subsequently, double intersection of any two variables $V=X\cap Y$ must be equal in size to the mutual information between them
$H(V)=I(X;Y)$ (109)
At the same time it is naturally assumed that the intersection is a subset of both $X$ and $Y$. Here we show that it is not always possible to find such $V$ using a simple example of binary random variables.
Lemma 1.
A mutual subvariable $V$ of two random variables $X$ and $Y$ such that its entropy is equal to the mutual information $I(X;Y)$ does not always exist
$\exists X,Y:\forall V\subset X,Y\hookrightarrow H(V)<I(X;Y)$ (110)
Proof.
Consider the following correlated fair coins
$X:\vec{p}(x)=\left(\begin{matrix}0.5\cr 0.5\end{matrix}\right),Y:\vec{p}(y)=\left(\begin{matrix}0.5\cr 0.5\end{matrix}\right),$ (111)
$p(x,y)=\left(\begin{matrix}0.1&0.4\cr 0.4&0.1\end{matrix}\right)$ (112)
Due to them being binary, a subvariable of each is either itself or a deterministic variable (i.e. empty set). They are clearly not equal, hence the only possibility is
$V=\emptyset\Rightarrow H(V)=0$ (113)
The mutual information $I(X;Y)$ is, however, non-zero. ∎
The operations of difference also has some peculiar properties. It can be shown to be not-unique. Consider the XOR example from the main text
$O_{1,2,3}=\begin{cases}0,&50\%\\ 1,&50\%\end{cases},$ (114)
$O_{3}=O_{1}\oplus O_{2}$ (115)
Taking a variable that represents the whole system $W=O_{1}\cup O_{2}\cup O_{3}$, we have two candidates $O_{2},O_{3}$ for the result of difference $W\backslash O_{1}$. Substituting them into the definition, we find out that both are valid, despite being explicitly unequal.
$O_{i}\cap O_{1}=\emptyset,$ (116)
$O_{i}\cup(W\cap O_{1})=O_{i}\cup O_{1}=W$ (117)
While distributivity axiom is broken in the space of random variables, a weaker condition ought to hold in any space ’similar’ to set theory.
Lemma 2.
As long as the relation of subset is well-defined in the following way
$X\subset Y\Leftrightarrow X\cup Y=Y\Leftrightarrow X\cap Y=X$ (118)
the subdistributivity between union and intersection holds
$(X\cap Z)\cup(Y\cap Z)\subset(X\cup Y)\cap Z$ (119)
Proof.
Let us prove first that lemma’s condition implies for any variables $X,Y,Z$
$X\subset Z,Y\subset Z\Rightarrow(X\cup Y)\subset Z$ (120)
Indeed, $(X\cup Y)\cup Z=(X\cup Z)\cup Y=Z\cup Y=Z$. In relation to our setup it is easy to see that
$(X\cap Z)\subset(X\cup Y)\cap Z,$ (121)
$(X\cap Z)\cap((X\cup Y)\cap Z)=(X\cup Y)\cap(Z\cap(X\cap Z))=(X\cup Y)\cap(X\cap Z)=(X\cap Z)$ (122)
Likewise $(Y\cap Z)\subset(X\cup Y)\cap Z$. By applying relation (120), we get the statement of the lemma
$(X\cap Z)\cup(Y\cap Z)\subset(X\cup Y)\cap Z$ (123)
Corollary.
Vice-versa, the relation $X\cup Y=Y\Leftrightarrow X\cap Y=X$ follows from subdistributivity.
To overcome the constraint of the random variable space being insufficient for the purpose of describing general information intersections, we extend it by adding all necessary elements to make set-theoretic operations well-defined. The extended RV space is a space with operations of union $\cup$ and intersection $\cap$ obeying a set of axioms:
$\forall X,Y,Z$ (124)
$X\cup\emptyset$ $=X$ $X\cap\emptyset$ =$\emptyset$= ∅
$X\cup X$ $=X$ $X\cap X$ $=X$
$X\cup Y$ $=Y\cup X$ $X\cap Y$ $=Y\cap X$
$(X\cup Y)\cup Z$ $=X\cup(Y\cup Z)$ $(X\cap Y)\cap Z$ $=X\cap(Y\cap Z)$
The relation of subsets is defined as follows
$X\subset Y\Leftrightarrow X\cap Y=X\Leftrightarrow X\cup Y=Y$ (125)
The difference of two members $X,Y$ is any $D=X\backslash Y$ such that
$D\cap Y=\emptyset,$ (126)
$D\cup(X\cap Y)=X$ (127)
Note that the result of taking the difference is not always unique (116).
After extending the random variables space, we prove that the inclusion-exclusion principle for $2$ variables, as expected, remains unmodified by the lack of distributivity
Lemma 3.
The size of the union of two extended RV space members is related to their own sizes and the size of their intersection:
$\hat{H}(X\cup Y)=\hat{H}(X)+\hat{H}(Y)-\hat{H}(X\cap Y)$ (128)
Proof.
First, we will rewrite the left side $H(X\cup Y)$ as a union of two disjoint pieces. Using the fact that difference $Y\backslash X$ is a subset of $Y$, we get
$X\cup(Y\backslash X)=X\cup Y$ (129)
At the same time, by definition $X\cap(Y\backslash X)=\emptyset$. It means that we may use the second axiom of the measure to write the size of union as a sum of sizes of its disjoint parts
$\hat{H}(X\cup Y)=\hat{H}(X)+\hat{H}(Y\backslash X)$ (130)
We may repeat the same steps in order to decompose the second summand into two more terms and obtain the statement of the lemma.
$(Y\backslash X)\cap(X\cap Y)=((Y\backslash X)\cap X)\cap Y=\emptyset,$ (131)
$(Y\backslash X)\cup(X\cap Y)=Y,$ (132)
$\hat{H}(Y)=\hat{H}(Y\backslash X)+\hat{H}(X\cap Y)$ (133)
The inclusion-exclusion principle for $3$ variables, however, along with all the terms from the set-theoretic expression, receives a peculiar extra term related to the breaking of distributivity
Theorem 1.
The size of triple union is related to the sizes of individual terms, their intersections and includes an additional term that measures the breaking of distributivity:
$\hat{H}(X_{1}\cup X_{2}\cup X_{3})=$ (134)
$=\hat{H}(X_{1})+\hat{H}(X_{2})+\hat{H}(X_{3})-\hat{H}(X_{1}\cap X_{2})-\hat{H}(X_{1}\cap X_{3})-\hat{H}(X_{2}\cap X_{3})+\hat{H}(X_{1}\cap X_{2}\cap X_{3})-\Delta\hat{H},$ (135)
where the last term is symmetric with respect to permutations of indices and equal to
$\Delta\hat{H}=\hat{H}(((X_{i}\cup X_{j})\cap X_{k})\backslash((X_{i}\cap X_{k})\cup(X_{j}\cap X_{k})))$ (136)
$\forall i\neq j\neq k$ (137)
Proof.
We begin by choosing 2 of 3 extended RV members on the left side and grouping them in order to use the result of the previous lemma (128).
$\hat{H}((X_{1}\cup X_{2})\cup X_{3})=\hat{H}(X_{1}\cup X_{2})+\hat{H}(X_{3})-\hat{H}((X_{1}\cup X_{2})\cap X_{3})$ (138)
First 2 terms are easily decomposed further, however for the last one the subdistributivity axiom has to be applied. Defining the breaking of distributivity as
$\Delta X_{123}=((X_{1}\cup X_{2})\cap X_{3})\backslash((X_{1}\cap X_{3})\cup(X_{2}\cap X_{3})),$ (139)
we can present the term as a disjoint union, which is then substituted into the second axiom of measure to give a sum
$\hat{H}((X_{1}\cup X_{2})\cap X_{3})=\hat{H}(((X_{1}\cap X_{3})\cup(X_{2}\cap X_{3}))\cup\Delta X_{123})=\hat{H}((X_{1}\cap X_{3})\cup(X_{2}\cap X_{3}))+\hat{H}(\Delta X_{123})$ (140)
This is the point where subdistributivity plays a crucial role. The first equality is made possible by the fact that the union of intersections is indeed a subset of the left side. Physically, as it will be clear later, it guaranties the non-negativity of synergy in the PID. Applying (128) once again, we get
$\hat{H}((X_{1}\cap X_{3})\cup(X_{2}\cap X_{3}))=\hat{H}(X_{1}\cap X_{3})+\hat{H}(X_{2}\cap X_{3})-\hat{H}(X_{1}\cap X_{2}\cap X_{3})$ (141)
We decompose $\hat{H}(X_{1}\cup X_{2})$ and combine everything into the final form
$\hat{H}(X_{1}\cup X_{2}\cup X_{3})=$ (142)
$=\hat{H}(X_{1})+\hat{H}(X_{2})+\hat{H}(X_{3})-\hat{H}(X_{1}\cap X_{2})-\hat{H}(X_{1}\cap X_{3})-\hat{H}(X_{2}\cap X_{3})+\hat{H}(X_{1}\cap X_{2}\cap X_{3})-\hat{H}(\Delta X_{123})$ (143)
The only term that depends on the order of putting brackets in (138) is $\hat{H}(\Delta X_{123})$. Due to the associativity and commutativity of both union and intersection, we conclude that while the quantity measured by the last term is explicitly not symmetric wrt permutations of indices
$\Delta X_{123}\neq\Delta X_{132}\neq\Delta X_{231},$ (144)
its size is always the same. Defining a single function equal to this size concludes the proof.
$\Delta\hat{H}=\hat{H}(\Delta X_{123})$ (145)
Appendix B Antichains
The antichain notation proved very useful in the PID framework [6, 11]. Its goal is to allow a simpler description for various information pieces. Each joint distribution is denoted by the collection of variables’ indices
$(X_{i_{1}},X_{i_{2}},\dots,X_{i_{m}})\rightarrow\{i_{1}i_{2}\dots i_{m}\}=\mathbf{A}$ (146)
There is a trivial partial order on such collections, based on the inclusion order of the corresponding distributions
$\mathbf{A}\subset\mathbf{B}\Leftrightarrow\forall i\in\mathbf{A}\hookrightarrow i\in\mathbf{B}$ (147)
To represent the intersections, a set of strong antichains is taken on the above poset
$\alpha=\mathbf{A}_{1}\mathbf{A}_{2}\dots\mathbf{A}_{n}=\{i_{11}\dots i_{1m_{1}}\}\{i_{21}\dots i_{2m_{2}}\}\dots\{i_{n1}\dots i_{nm_{n}}\},$ (148)
$\forall a,b,c,d\hookrightarrow i_{ab}\neq i_{cd}$ (149)
It forms a new lattice $(L^{\cap},\preceq)$, which is referred to as the redundancy lattice
$\forall\alpha,\beta\in L^{\cap}$ (150)
$\alpha\preceq\beta\Leftrightarrow\forall\mathbf{B}\in\beta\text{ }\exists\mathbf{A}\in\alpha:\mathbf{A}\preceq\mathbf{B}$ (151)
This lattice turns out to be isomorphic to the information lattice $(L^{I},\preceq)$ from the partitioning algorithm via the mapping
$\forall\alpha=\{i_{11}\dots i_{1m_{1}}\}\{i_{21}\dots i_{2m_{2}}\}\dots\{i_{n1}\dots i_{nm_{n}}\},$ (152)
$\alpha\leftrightarrow\Xi_{\alpha}=\bigcap_{j=\overline{1,n}}\left(\bigcup_{k=\overline{1,m_{j}}}X_{i_{jk}}\right)$ (153)
In order to prove this [theorem 2], we need several additional lemmas
Lemma 4.
For any two antichains such that
$\alpha=\mathbf{A}\mathbf{A}_{1}\mathbf{A}_{2}\dots,$ (154)
$\beta=\mathbf{B}\mathbf{A}_{1}\mathbf{A}_{2}\dots,$ (155)
$\mathbf{B}\preceq\mathbf{A}$ (156)
the corresponding partitioning algorithm terms are comparable as
$\Xi_{\beta}\preceq\Xi_{\alpha}$ (157)
Proof.
Without loss of generality, let the differing index collections be
$\mathbf{A}=\{12\dots N\},$ (158)
$\mathbf{B}=\{12\dots n\},$ (159)
$n<N$ (160)
Now consider a partitioning algorithm step
$\Xi_{\alpha}=(X_{1}\cup X_{2}\cup\dots X_{N})\cap\Xi^{\prime},$ (161)
$\Xi_{\beta}=(X_{1}\cup X_{2}\cup\dots X_{n})\cap\Xi^{\prime},$ (162)
$\Xi_{\gamma}=(X_{n+1}\cup X_{2}\cup\dots X_{N})\cap\Xi^{\prime},$ (163)
$\Xi_{\delta}=(X_{1}\cup X_{2}\cup\dots X_{n})\cap(X_{n+1}\cup X_{2}\cup\dots X_{N})\cap\Xi^{\prime}$ (164)
By the definition of the information lattice, we have
$\Xi_{\beta}\subset\Xi_{\alpha}\Rightarrow\Xi_{\beta}\preceq\Xi_{\alpha}$ (165)
Lemma 5.
For any two antichains such that
$\gamma=\mathbf{A}_{1}\mathbf{A}_{2}\dots,$ (166)
$\delta=\mathbf{A}\mathbf{A}_{1}\mathbf{A}_{2}\dots,$ (167)
the corresponding partitioning algorithm terms are comparable as
$\Xi_{\delta}\preceq\Xi_{\gamma}$ (168)
Proof.
Without loss of generality, we shall put the indices of the first two collections to be
$\mathbf{A}=\{12\dots n\},$ (169)
$\mathbf{A}_{1}=\{(n+1)\dots N\},$ (170)
$n<N$ (171)
Now consider a partitioning algorithm step
$\Xi_{\alpha}=(X_{1}\cup X_{2}\cup\dots X_{N})\cap\Xi^{\prime},$ (172)
$\Xi_{\beta}=(X_{1}\cup X_{2}\cup\dots X_{n})\cap\Xi^{\prime},$ (173)
$\Xi_{\gamma}=(X_{n+1}\cup X_{2}\cup\dots X_{N})\cap\Xi^{\prime},$ (174)
$\Xi_{\delta}=(X_{1}\cup X_{2}\cup\dots X_{n})\cap(X_{n+1}\cup X_{2}\cup\dots X_{N})\cap\Xi^{\prime}$ (175)
By the definition of the information lattice, we have
$\Xi_{\delta}\subset\Xi_{\gamma}\Rightarrow\Xi_{\delta}\preceq\Xi_{\gamma}$ (176)
$(X_{1}\cup X_{2}\cup\dots\cup X_{m})\cap\Xi^{\prime}$$(X_{1}\cup\dots\cup X_{i})\cap\Xi^{\prime}$$(X_{i+1}\cup\dots\cup X_{m})\cap\Xi^{\prime}$$(X_{1}\cup\dots\cup X_{i})\cap(X_{i+1}\cup\dots\cup X_{m})\cap\Xi^{\prime}$
Figure 6: Information lattice piece obtained from a single partitioning step. The arrows represent the inclusion order. The pieces from all possible steps and partitions are then combined into a full information lattice simply by identifying equal regions $\Xi$.
Theorem 2.
The information lattice $(L^{I},\preceq)$ is isomorphic to the lattice of antichains $(L^{\cap},\preceq)$. The canonical isomorphism is given by (152).
Proof.
$\Leftarrow$
The information lattice is built from the pieces appearing during the partitioning algorithm of the form [Figure 6]. Using the notation (152) we prove that for the corresponding antichains the same relations hold
$\Xi_{\alpha}=(X_{1}\cup X_{2}\cup\dots\cup X_{m})\cap\Xi^{\prime},$ (177)
$\Xi_{\beta}=(X_{1}\cup\dots\cup X_{i})\cap\Xi^{\prime},$ (178)
$\Xi_{\gamma}=(X_{i+1}\cup\dots\cup X_{m})\cap\Xi^{\prime},$ (179)
$\Xi_{\delta}=\left((X_{1}\cup\dots\cup X_{i})\cap\Xi^{\prime}\right)\cap\left((X_{i+1}\cup\dots\cup X_{m})\cap\Xi^{\prime}\right),$ (180)
$\alpha=\{1\dots m\}\dots,$ (181)
$\beta=\{1\dots i\}\dots,$ (182)
$\gamma=\{i+1\dots m\}\dots,$ (183)
$\delta=\{1\dots i\}\{i+1\dots m\}\dots$ (184)
$\beta\preceq\alpha,$ (185)
$\gamma\preceq\alpha,$ (186)
$\delta\preceq\beta,$ (187)
$\delta\preceq\gamma,$ (188)
The lattice $(L^{I},\preceq)$ follows from $(L^{\cap},\preceq)$.
$\Rightarrow$
For any two antichains $\beta\preceq\alpha$, by definition we have without loss of generality
$\alpha=\mathbf{A}_{1}\mathbf{A}_{2}\dots\mathbf{A}_{n},$ (189)
$\beta=\mathbf{B}_{1}\mathbf{B}_{2}\dots\mathbf{B}_{N},$ (190)
$\beta\preceq\alpha\Rightarrow n<N,\forall i=\overline{1,n}\hookrightarrow\mathbf{B}_{i}\preceq\mathbf{A}_{i}$ (191)
By lemma 4, for the set of antichains $\alpha_{i}=\mathbf{B}_{1}\mathbf{B}_{2}\dots\mathbf{B}_{i}\mathbf{A}_{i+1}\dots\mathbf{A}_{n}$, we get
$\Xi_{\alpha_{n}}\preceq\Xi_{\alpha_{n-1}}\preceq\dots\preceq\Xi_{\alpha_{1}}\preceq\Xi_{\alpha}$ (192)
By lemma 5, for the set of antichains $\beta_{i}=\mathbf{B}_{1}\mathbf{B}_{2}\dots\mathbf{B}_{i}$, we get
$\Xi_{\beta}\preceq\Xi_{\beta_{N-1}}\preceq\dots\preceq\Xi_{\beta_{n+1}}\preceq\Xi_{\alpha_{n}}$ (193)
This proves
$\Xi_{\beta}\preceq\Xi_{\alpha},$ (194)
i.e. lattice $(L^{\cap},\preceq)$ follows from $(L^{I},\preceq)$. ∎
In the next section we will be heavily using this equivalence by denoting the partitioning terms $\Xi$ using antichains. It is useful to note that the covering $C$ of the partition term $\Xi_{\alpha}$ is always equal to the cardinality of the corresponding antichain
$C=|\alpha|$ (195)
Appendix C Information atoms
The first index of the parthood table $f_{ij}$ denotes the partitioning term $\Xi_{i}$. Using the results of the previous section, we may replace it with an antichain $\alpha$
$\Xi_{\alpha j}=\sum_{j}f_{\alpha j}\Pi_{j},$ (196)
$C_{\alpha}=|\alpha|$ (197)
With this in mind, the axioms of the decomposition may be rewritten in a simpler form
$\alpha\preceq\beta\Rightarrow\forall i\hookrightarrow f_{\alpha i}\leq f_{\beta i},$ (198)
$c_{i}=\max_{\alpha:f_{\alpha i}=1}|\alpha|$ (199)
$\sum_{k=1}^{N}H(X_{k})=\sum_{i}c_{i}\Pi_{i}$ (200)
Several relations on the information atoms derived for a general setup will be useful in the following discussion. The first lemma proves that one cannot fit new atoms in the information lattice node, whose decomposition is already fully fixed by the lower nodes.
Lemma 6.
If two partitioning terms are equal in size and comparable in the information lattice, they must have the same decomposition.
$\hat{H}(\Xi_{\alpha})=\hat{H}(\Xi_{\beta}),\alpha\preceq\beta\Rightarrow\forall i\text{ }f_{\alpha i}=f_{\beta i}$ (201)
Proof.
By monotonicity $\forall i$
$f_{\alpha i}=1\Rightarrow f_{\beta i}=1$ (202)
Hence, to be different, $f_{\beta i}$ must include at least one extra atom
$\exists j:f_{\alpha j}\neq f_{\beta j}\Rightarrow f_{\alpha j}=0,f_{\beta j}=1$ (203)
We immediately get a contradiction with the non-negativity of information, since existing atoms already occupy the whole size of $\Xi_{\beta}$
$\hat{H}(\Xi_{\beta})-\hat{H}(\Xi_{\alpha})=0=\Pi_{j}+\dots\geq\Pi_{j}\Rightarrow\Pi_{j}=0$ (204)
The second lemma states that for two explicitly equal partition terms the result of the decomposition is also the same.
Lemma 7.
$\Xi_{\alpha}=\Xi_{\beta}\Rightarrow\forall i\text{ }f_{\alpha i}=f_{\beta i}$ (205)
Proof.
For the information lattice node $\Xi_{\gamma}=\Xi_{\alpha}\cap\Xi_{\beta}$ it is true that
$\gamma\preceq\alpha,\gamma\preceq\beta,$ (206)
$\hat{H}(\Xi_{\gamma})=\hat{H}(\Xi_{\alpha})=\hat{H}(\Xi_{\beta})$ (207)
hence, by lemma 6
$\forall i\text{ }f_{\gamma i}=f_{\alpha i}=f_{\beta i}$ (208)
Now we will use the linear structure of the solution space to prove that the superposition of two solutions for systems with equal number of variables $N$ is a solution to the superposition of these systems.
Lemma 8.
For any two sets of random variables $\{X_{i}\}_{i=\overline{1,N}}$ and $\{Y_{j}\}_{j=\overline{1,N}}$ let their information decompositions be correspondingly $\{\Pi_{i}(X_{1};\dots X_{N})\}$ and $\{\Pi_{i}(Y_{1};\dots Y_{N})\}$. Then, for a system $Z_{1},\dots Z_{N}$ such that for all joint distributions
$H(Z_{i_{1}},Z_{i_{2}},\dots Z_{i_{n}})=aH(X_{i_{1}},X_{i_{2}},\dots X_{i_{n}})+bH(Y_{i_{1}},Y_{i_{2}},\dots Y_{i_{n}})$ (209)
a possible solution is a superposition of two previous decompositions
$\forall i\hookrightarrow\Pi_{i}(Z_{1};\dots Z_{N})=a\Pi_{i}(X_{1};\dots X_{N})+b\Pi_{i}(Y_{1};\dots Y_{N}),$ (210)
Proof.
The formulas of the partitioning algorithm are linear, hence we may assume all $\Xi$ terms of the new system to be a superposition of the same terms from its constituents
$\forall\alpha\hookrightarrow\hat{H}(\Xi_{\alpha}(Z_{1};\dots Z_{N}))=a\hat{H}(\Xi_{\alpha}(X_{1};\dots X_{N}))+b\hat{H}(\Xi_{\alpha}(Y_{1};\dots Y_{N}))$ (211)
Defining the solution of the new system as
$\forall i\hookrightarrow\Pi_{i}(Z_{1};\dots Z_{N})=a\Pi_{i}(X_{1};\dots X_{N})+b\Pi_{i}(Y_{1};\dots Y_{N}),$ (212)
We indeed see that it complies with the parthood table
$\hat{H}(\Xi_{\alpha}(Z_{1};\dots Z_{N}))=\sum_{i}f_{\alpha i}\Pi_{k}(Z_{1};\dots Z_{N})=\sum_{i}f_{\alpha i}a\Pi_{i}(X_{1};\dots X_{N})+\sum_{i}f_{\alpha i}b\Pi_{i}(Y_{1};\dots Y_{N})=$ (213)
$=a\hat{H}(\Xi_{\alpha}(X_{1};\dots X_{N}))+b\hat{H}(\Xi_{\alpha}(Y_{1};\dots Y_{N}))$ (214)
The atom sizes changed, hence we need to check the solution against the information conservation law
$\sum_{i=1}^{N}H(Z_{i})=a\sum_{i=1}^{N}H(X_{i})+b\sum_{i=1}^{N}H(Y_{i})=\sum_{j}c_{j}a\Pi_{j}[c_{j}]+\sum_{j}c_{j}b\Pi_{j}[c_{j}]=\sum_{k}c_{k}(a+b)\Pi_{k}[c_{k}]$ (215)
The proved additivity of solutions opens the way to recover decompositions arbitrary systems by superposing more basic systems. We proceed to explore various degenerate systems with exact solutions.
Set-theoretic prediction
Let us start with the direct analogy of the Venn diagrams for sets - systems for which we demand full compliance with distributivity axiom. There is no variability in the inclusion-exclusion formulas and the atoms are recovered via classical Möbius inversion. To perform it, we use the same antichain lattice as the notation for atoms
$\Xi_{\alpha}=\sum_{\beta\preceq\alpha}\Pi_{\beta}[|\beta|]$ (216)
The collection of atoms is simplified even further, since the subset of $\Xi$-terms corresponding to antichains of the form $\iota=\{i_{1}\}\{i_{2}\}\dots\{i_{n}\}$ is already sufficient to define the whole system (all other equations become linearly dependent). Therefore, all non-zero atoms are indexed by $\{\iota\}$. Their sizes are recovered using the interaction information functions
$\Pi_{\iota}=\sum_{m=n}^{N}(-1)^{m-n}\sum_{i_{n+1},\dots i_{m}}I_{m}(X_{i_{1}};\dots X_{i_{m}}),$ (217)
where
$I_{N}(X_{1};\dots X_{N})\equiv\sum_{n=1}^{N}(-1)^{n-1}\sum_{i_{1},\dots i_{n}}H(X_{i_{1}},X_{i_{2}},\dots,X_{i_{n}})$ (218)
This also gives the validity criterion for such decompositions, since all the atoms must be non-negative
$\forall\{i_{1}\}\{i_{2}\}\dots\{i_{n}\}\hookrightarrow\sum_{m=n}^{N}(-1)^{m-n}\sum_{i_{n+1},\dots i_{m}}I_{m}(X_{i_{1}};\dots X_{i_{m}})\geq 0$ (219)
We can prove that for $N=3$ variables it condenses down to the $3^{\text{rd}}$ order interaction information being non-negative.
Lemma 9.
Any system of $3$ variables $X_{1},X_{2},X_{3}$ with $3^{\text{rd}}$-order interaction information function $I_{3}(X_{1};X_{2};X_{3})\geq 0$, may be decomposed using only non-negative set-theoretic atoms.
Proof.
$\forall i\neq j\neq k,$ (220)
$\Pi_{\{1\}\{2\}\{3\}}=I_{3}(X_{1};X_{2};X_{3})\geq 0,$ (221)
$\Pi_{\{i\}\{j\}}=I(X_{i};X_{j})-I_{3}(X_{1};X_{2};X_{3})=I(X_{i};X_{j}|X_{k})\geq 0,$ (222)
$\Pi_{\{i\}}=H(X_{i})-I(X_{i};X_{j})-I(X_{i};X_{k})+I_{3}(X_{1};X_{2};X_{3})=$ (223)
$=H(X_{i})-\left(2H(X_{i})+H(X_{j})+H(X_{k})-H(X_{i},X_{j})-H(X_{i},X_{k})\right)+$ (224)
$+\left(H(X_{i})+H(X_{j})+H(X_{k})-H(X_{i},X_{j})-H(X_{j},X_{k})-H(X_{i},X_{k})+H(X_{i},X_{j},X_{k})\right)=$ (225)
$=-H(X_{j},X_{k})+H(X_{i},X_{j},X_{k})\geq 0$ (226)
Set-theoretic systems never exhibit synergistic properties. The following result can be understood in the sense that the breaking of distributivity is a necessary condition for the existence of synergy in any $N$-variable system
Lemma 10.
In set-theoretic system mutual information is always subadditive
$I(X_{1};X_{N+1})+I(X_{2};X_{N+1})+\dots+I(X_{N};X_{N+1})\geq I((X_{1},X_{2},\dots X_{N});X_{N+1})$ (227)
Proof.
Let us substitute the atoms of decomposition (217) into the inequality.
$\sum_{k=1}^{N}\sum_{\iota^{\prime}\preceq\{k\}\{N+1\}}\Pi_{\iota}\geq\sum_{\iota\preceq\{12\dots N\}\{N+1\}}\Pi_{\iota}$ (228)
For any atom on the right side, we have by (150
$\iota\preceq\{12\dots N\}\{N+1\}\Rightarrow\exists\{i_{a}\},\{i_{b}\}\in\iota:\begin{cases}\{i_{a}\}\preceq\{12\dots N\}\\ \{i_{b}\}\preceq\{N+1\}\end{cases}$ (229)
Since $\iota$ is composed of single indices, we have
$i_{a}=\overline{1,N},$ (230)
$i_{b}=N+1$ (231)
Then such term can also be found on the left side of (228), as
$\iota\preceq\{i_{a}\}\{N+1\}$ (232)
The non-negativity of all atoms proves the statement of the lemma. ∎
XOR gate
The XOR gate [114] is a paradigmatic example of a synergistic system and a manifestation of non-set-theoretic properties of information. It contains a completely different set of atoms. With three mutually independent initial variables, the set of non-empty partition terms simplifies to
$\forall i\neq j\neq k$
$\Xi_{\{123\}}[1]$ $=O_{1}\cup O_{2}\cup O_{3}$ $\hat{H}(\Xi_{\{123\}})$ $=2,$
$\Xi_{\{ij\}}[1]$ $=O_{i}\cup O_{j}$ $\hat{H}(\Xi_{\{ij\}})$ $=2,$
$\Xi_{\{i\}}[1]$ $=O_{i}$ $\hat{H}(\Xi_{\{i\}})$ $=1,$
$\Xi_{\{ij\}\{k\}}[2]$ $=(O_{i}\cup O_{j})\cap O_{k}$ $\hat{H}(\Xi_{\{ij\}\{k\}})$ $=1$
$\Xi_{\{123\}}$$\Xi_{\{13\}}$$\Xi_{\{12\}}$$\Xi_{\{23\}}$$\Xi_{\{1\}}$$\Xi_{\{2\}}$$\Xi_{\{3\}}$$\Xi_{\{12\}\{3\}}$$\Xi_{\{13\}\{2\}}$$\Xi_{\{23\}\{1\}}$
Figure 7: Information lattice for the XOR gate example. Empty nodes are omitted.
It is possible to guess the solution, starting from the lower nodes in the lattice [Fig. 7] and introducing more atoms for each higher node. By lemma 6, new atoms appear only in $\Xi_{\{ij\}\{k\}}$ and $\Xi_{\{ij\}}$. Due to the symmetry of the problem, there are two variants for the decomposition of the former: $3$ distinct atoms $\Pi_{\{ij\}\{k\}}$ or one symmetrically shared $\Pi_{s}$. The coverings are found from (198).
$\hat{H}(\Xi_{\{ij\}\{k\}}[2])=\hat{H}(\Xi_{\{k\}}[1])=\Pi_{\{ij\}\{k\}}[2]+\Pi_{s}[2]$ (233)
Since the union of any two XOR variables defines the whole system, we may use lemma 7. The parthood table columns are thus the same for $\{12\}$, $\{13\}$, $\{23\}$ and $\{123\}$. A single shared atom $\Pi_{g}$ may be added in the most general case.
$\hat{H}(\Xi_{\{ij\}}[1])=\hat{H}(\Xi_{\{123\}}[1])=\Pi_{\{12\}\{3\}}[2]+\Pi_{\{13\}\{2\}}[2]+\Pi_{\{23\}\{1\}}[2]+\Pi_{s}[2]+\Pi_{g}[1]$ (234)
A solution is then obtained from the information conservation law with only two non-zero atoms
$\Pi_{s}[2]=1,$ (235)
$\Pi_{g}[1]=1$ (236)
It is a unique symmetric solution for the XOR gate:
Theorem 3.
The XOR gate has a unique symmetric solution (Table 1)
Proof.
Assuming a general form of the parthood table, we may add several more atoms compared to our previous discussion:
$\hat{H}(\Xi_{\{ij\}\{k\}}[2])=\Pi_{{\{ij\}\{k\}}}[2]+\Pi_{\{ij\}\{k\},\{ik\}\{j\}}[2]+\Pi_{\{ij\}\{k\},\{kj\}\{i\}}[2]+\Pi_{s}[2],$ (237)
$f$$\{1\}\{2\}\{3\}$$\{1\}\{2\}$$\{1\}\{3\}$$\{2\}\{3\}$$\{12\}\{3\}$$\{13\}\{2\}$$\{23\}\{1\}$$\{1\}$$\{2\}$$\{3\}$$\{12\}$$\{13\}$$\{23\}$$\{123\}$
$\Pi_{s}$ 0 0 0 0 1 1 1 1 1 1 1 1 1 1
$\Pi_{g}$ 0 0 0 0 0 0 0 0 0 0 1 1 1 1
$\Pi_{{\{12\}\{3\}}}$ 0 0 0 0 1 0 0 0 0 1 1 1 1 1
$\Pi_{{\{13\}\{2\}}}$ 0 0 0 0 0 1 0 0 1 0 1 1 1 1
$\Pi_{{\{23\}\{1\}}}$ 0 0 0 0 0 0 1 1 0 0 1 1 1 1
$\Pi_{{\{12\}\{3\}},{\{13\}\{2\}}}$ 0 0 0 0 1 1 0 0 1 1 1 1 1 1
$\Pi_{{\{13\}\{2\},{\{23\}\{1\}}}}$ 0 0 0 0 0 1 1 1 1 0 1 1 1 1
$\Pi_{{\{23\}\{1\},{\{12\}\{3\}}}}$ 0 0 0 0 1 0 1 1 0 1 1 1 1 1
In a symmetric solution, they are all equal, hence let
$\Pi_{s}=x,$ (238)
$\Pi_{\{ij\}\{k\},\{kj\}\{i\}}=y,$ (239)
$\Pi_{{\{ij\}\{k\}}}=1-2y-x,$ (240)
$\Pi_{g}=2-x-3y-3(1-2y-x)=2x+3y-1$ (241)
The conservation law now dictates that
$\sum_{i}H(X_{i})=3=2x+2*3y+2*3*(1-2y-x)+2x+3y-1=5-2x-3y,$ (242)
$2\displaystyle 2x+3y=$ 2 (243)
But we know that all atoms have non-negative sizes, hence
$\Pi_{{\{ij\}\{k\}}}=1-2y-x=-0.5y\geq 0,$ (244)
$y=0,$ (245)
$x=1$ (246)
And we get the same solution
$\Pi_{s}=1,$ (247)
$\Pi_{g}=1$ (248)
Table 1: Parthood table for the unique XOR gate solution
$f$$\{1\}\{2\}\{3\}$$\{1\}\{2\}$$\{1\}\{3\}$$\{2\}\{3\}$$\{12\}\{3\}$$\{13\}\{2\}$$\{23\}\{1\}$$\{1\}$$\{2\}$$\{3\}$$\{12\}$$\{13\}$$\{23\}$$\{123\}$
$\Pi_{s}$ 0 0 0 0 1 1 1 1 1 1 1 1 1 1
$\Pi_{g}$ 0 0 0 0 0 0 0 0 0 0 1 1 1 1
The corresponding parthood table may be found in (Table 1). From it we can see that while the symmetric atom $\Pi_{s}$ appears in each variable $X_{i},i=\overline{1,3}$, it is not a part of any intersection $X_{i}\cap X_{j}$. Using the XOR gate example, we may show that information atoms cannot be described as members of the extended RV space
Lemma 11.
For a system of random variables $X_{1},\dots X_{N}$ there does not always exist a universal collection of disjoint extended RV space members with covering numbers ${\it\Pi}_{j}[c_{j}]$, such that all partition terms are constructed from them
$\Xi_{\alpha}=\bigcup_{j}{\it\Pi}_{j},$ (249)
they comply with the covering numbers of partition terms
${\it\Pi}_{j}[c_{j}]\subset\Xi_{\alpha}\Rightarrow c_{j}\geq|\alpha|$ (250)
and satisfy the information conservation law
$\sum_{i=1}^{N}H(X_{i})=\sum_{j}c_{j}\hat{H}({\it\Pi}_{j}[c_{j}])$ (251)
Proof.
First, let us show that in extended RV if $C\subset A,C\subset B$, then $C\subset(A\cap B)$
$C\cap(A\cap B)=(C\cap A)\cap B=C\cap B=C$ (252)
Now assume that we have found an invariant collection ${\it\Pi}_{j}$ for the XOR setup. For each variable we have
$H(O_{1})=\sum_{j_{1}}\hat{H}({\it\Pi}_{j_{1}}),$ (253)
$H(O_{2})=\sum_{j_{2}}\hat{H}({\it\Pi}_{j_{2}}),$ (254)
$H(O_{3})=\sum_{j_{3}}\hat{H}({\it\Pi}_{j_{3}}),$ (255)
where the sets of atoms $j_{1},j_{2},j_{3}$ do not intersect due to our observation (252) and since the variables are independent $O_{i_{1}}\cap O_{i_{2}\neq i_{1}}=\emptyset$.
For each of 3 possible partitions there is a single region $\{i_{1}i_{2}\}\{i_{3}\}$ with the covering by definition $C=2$. One can notice that
$\forall i_{1}\neq i_{2}\neq i_{3}\hookrightarrow(O_{i_{1}}\cup O_{i_{2}})\cap O_{i_{3}}=O_{i_{3}}$ (256)
Hence, for any $j_{a=\overline{1,3}}$, ${\it\Pi}_{j_{a}}$ is covered at least twice $c_{j_{a}}\geq 2$. If the procedure is invariant to the choice of partition, this is true for all $a$ at once. This, of course, contradicts the information conservation law, since we essentially doubled the entire system
$3=\sum_{i}H(O_{i})=\sum_{a=1}^{3}\sum_{j_{a}}\hat{H}({\it\Pi}_{j_{a}}),$ (257)
$\sum_{\text{all atoms}}c_{j}\hat{H}({\it\Pi_{j}})\geq\sum_{a=1}^{3}\sum_{j_{a}}c_{j_{a}}\hat{H}({\it\Pi}_{j_{a}})\geq 2\sum_{a=1}^{3}\sum_{j_{a}}\hat{H}({\it\Pi}_{j_{a}})=6,$ (258)
$\sum_{j}H(O_{j})=3<6\leq\sum_{\text{all atoms}}c_{j}\hat{H}({\it\Pi_{j}})$ (259)
One can see that due to the 2-covered atom $\Pi_{s}$ ’jumping’ between the variables depending on the chosen partition, we may not assign specific covering numbers to the extended RV space regions in a universal way.
N-parity
A generalization of XOR gate is the $N$-parity setup, also a symmetric system, for which one of the variables is defined by the combination of all the others
$X_{\overline{1,N}}=\begin{cases}0,&50\%\\ 1,&50\%\end{cases},$ (260)
$\forall i=\overline{1,N}\hookrightarrow X_{i}\equiv\sum_{j\neq i}X_{j}\mod 2$ (261)
There are only two types of non-empty partition terms here. Firstly, there are simple $1$-covered joint distributions
$\Xi_{\{i_{1}i_{2}\dots i_{n}\}}[1]=\bigcup_{k=\overline{1,n}}X_{i_{k}},$ (262)
$\hat{H}(\Xi_{\{i_{1}i_{2}\dots i_{n}\}})=\min(n,N-1)$ (263)
Secondly, we may get an intersection of covering $C=2$, as any $N-1$ variables completely determine the last one.
$\Xi_{\{i_{1}\}\{i_{2}i_{3}\dots i_{N}\}}[1]=X_{i_{1}}\cap\left(\bigcup_{k=\overline{2,N}}X_{i_{k}}\right),$ (264)
$\hat{H}(\Xi_{\{i_{1}\}\{i_{2}i_{3}\dots i_{N}\}})=1$ (265)
Even more generally, dividing all $N$ variables into two sets and intersecting corresponding joint distributions gives us a $1$-bit $2$-covered term.
$\Xi_{\{i_{1}i_{2}\dots i_{n}\}\{i_{n+1}\dots i_{N}\}}[1]=\left(\bigcup_{k=\overline{1,n}}X_{i_{k}}\right)\cap\left(\bigcup_{l=\overline{n+1,N}}X_{i_{l}}\right),$ (266)
$\hat{H}(\Xi_{\{i_{1}i_{2}\dots i_{n}\}\{i_{n+1}\dots i_{N}\}})=1$ (267)
A solution can be easily guessed in a similar form to the XOR. Let there be a single symmetric $2$-covered atom $\Pi_{s}[2]=1$ and a set of $N-2$ ghost atoms $\Pi_{g_{k}}[1]=1,k=\overline{1,N-2}$, such that
$\Xi_{\{i_{1}i_{2}\dots i_{n}\}\{i_{n+1}\dots i_{N}\}}=\Pi_{s},$ (268)
$\Xi_{\{i_{1}i_{2}\dots i_{n}\}}=\Pi_{s}+\sum_{k=1}^{n-2}\Pi_{g_{k}}$ (269)
We immediately see that the information conservation law is satisfied
$\sum_{i}H(X_{i})=N=2\Pi_{s}+\sum_{k=1}^{N-2}\Pi_{g_{k}}$ (270)
N-parity with copies
One could get a more complex picture by adding multiple ’copies’ of each variable into the $N$-parity setup. Consider a set of variables
$X_{ij}:i=\overline{1,N};j=\overline{1,m_{i}};X_{ij}=\begin{cases}0,&50\%\\ 1,&50\%\end{cases}$ (271)
$\forall j=\overline{1,m_{i}}\hookrightarrow X_{ij}=X_{i1},$ (272)
$\forall i=\overline{1,N}\hookrightarrow X_{i1}\equiv\sum_{j\neq i}X_{j1}\mod 2,$ (273)
Without loss of generality, let us order the equivalence class sizes $m_{i}$
$m_{1}\geq m_{2}\geq\dots\geq m_{N}$ (274)
We will assume a solution of the same form as in the ordinary $N$-parity with the only difference being unknown covering numbers, which will depend on the set of $m_{i}$’s.
$\Pi_{s}[c_{s}]=1,$ (275)
$\forall n=\overline{1,N-2}\hookrightarrow\Pi_{g_{n}}[c_{g_{n}}]=1,$ (276)
Partitions yield several kinds of terms. Any intersection of two variables from different groups (i.e. with different first indices) is empty. Therefore, we may generalize the terms from previous setup (262) by taking $M$ copies of each variable in the joint distribution
$\Xi^{(1)}=\bigcap_{l={1,M}}\left(\bigcup_{k=\overline{1,n}}X_{i_{k}j_{l}}\right),$ (277)
$\hat{H}(\Xi^{(1)})=\min(n,N-1),$ (278)
$C=M\leq\min_{k=\overline{1,n}}(m_{k})$ (279)
The generalized term of the second kind (266) will have the following form
$\Xi^{(2)}=\left(\bigcap_{l={1,M}}\bigcup_{k=\overline{1,n}}X_{i_{k}j_{l}}\right)\cap\left(\bigcap_{l^{\prime}={1,M^{\prime}}}\bigcup_{k^{\prime}=\overline{n+1,N}}X_{i_{k^{\prime}}j_{l^{\prime}}}\right),$ (281)
$\hat{H}(\Xi^{(2)})=1,$ (282)
$C=M+M^{\prime}\leq m_{1}+m_{N}$ (283)
The maximal covering occurs in the case when one of the brackets has a single element $M=1$. The covering of the shared atom is then
$c_{s}=m_{1}+m_{N},$ (284)
$\hat{H}(\Xi^{(2)})=\Pi_{s}$ (285)
The ghost atoms $\Pi_{g_{k}}$ are found in the terms of first kind [277] with $k=n-1$. Thus they get coverings each equal to $(k+1)$-th maximal size $m_{k+1}$.
$\forall k=\overline{1,N-2}\hookrightarrow c_{g_{k}}=m_{k+1},$ (286)
$\hat{H}(\Xi^{(1)})=\Pi_{s}+\sum_{k=1}^{n-1}\Pi_{g_{k}}$ (287)
This fits exactly into the information conservation law
$\sum_{i,j}H(X_{ij})=\sum_{k=1}^{N}m_{k}=c_{s}+\sum_{n=1}^{N-2}c_{g_{n}}$ (288)
Arbitrary trivariate system
Combining the solutions of XOR gate with the set-theoretic prediction for $3$ variables using lemma 8, we get a solution for an arbitrary trivariate system (Table 2). The last columns of the parthood table may be written as a system of equations on atom sizes
$H(X_{1})=\Pi_{\{1\}\{2\}\{3\}}[3]+\Pi_{s}[2]+\Pi_{\{1\}\{2\}}[2]+\Pi_{\{1\}\{3\}}[2]+\Pi_{\{1\}}[1],$ (289)
$H(X_{2})=\Pi_{\{1\}\{2\}\{3\}}[3]+\Pi_{s}[2]+\Pi_{\{1\}\{2\}}[2]+\Pi_{\{2\}\{3\}}[2]+\Pi_{\{2\}}[1],$ (290)
$H(X_{3})=\Pi_{\{1\}\{2\}\{3\}}[3]+\Pi_{s}[2]+\Pi_{\{1\}\{3\}}[2]+\Pi_{\{2\}\{3\}}[2]+\Pi_{\{3\}}[1],$ (291)
$H(X_{1},X_{2})=\Pi_{\{1\}\{2\}\{3\}}[3]+\Pi_{s}[2]+\Pi_{\{1\}\{2\}}[2]+\Pi_{\{1\}\{3\}}[2]+\Pi_{\{2\}\{3\}}[2]+$ (292)
$+\Pi_{\{1\}}[1]+\Pi_{\{2\}}[1]+\Pi_{g}[1],$ (293)
$H(X_{1},X_{3})=\Pi_{\{1\}\{2\}\{3\}}[3]+\Pi_{s}[2]+\Pi_{\{1\}\{2\}}[2]+\Pi_{\{1\}\{3\}}[2]+\Pi_{\{2\}\{3\}}[2]+$ (294)
$+\Pi_{\{1\}}[1]+\Pi_{\{3\}}[1]+\Pi_{g}[1],$ (295)
$H(X_{2},X_{3})=\Pi_{\{1\}\{2\}\{3\}}[3]+\Pi_{s}[2]+\Pi_{\{1\}\{2\}}[2]+\Pi_{\{1\}\{3\}}[2]+\Pi_{\{2\}\{3\}}[2]+$ (296)
$+\Pi_{\{2\}}[1]+\Pi_{\{3\}}[1]+\Pi_{g}[1],$ (297)
$H(X_{1},X_{2},X_{3})=$ (298)
$=\Pi_{\{1\}\{2\}\{3\}}[3]+\Pi_{s}[2]+\Pi_{\{1\}\{2\}}[2]+\Pi_{\{1\}\{3\}}[2]+\Pi_{\{2\}\{3\}}[2]+$ (299)
$+\Pi_{\{1\}}[1]+\Pi_{\{2\}}[1]+\Pi_{\{3\}}[1]+\Pi_{g}[1],$ (300)
$\Pi_{s}[2]=\Pi_{g}[1]$ (301)
Lemma 12.
Any system of three random variable can be decomposed into a set of non-negative atoms (289).
Proof.
One can find the sizes of atoms $\Pi_{\{i\}}$ from the last $4$ equations in the system
$\Pi_{\{1\}}=H(X_{1},X_{2},X_{3})-H(X_{2},X_{3})\geq 0,$ (302)
$\Pi_{\{2\}}=H(X_{1},X_{2},X_{3})-H(X_{1},X_{3})\geq 0,$ (303)
$\Pi_{\{3\}}=H(X_{1},X_{2},X_{3})-H(X_{1},X_{2})\geq 0$ (304)
For the rest of set-theoretic atoms we have
$I(X_{1};X_{2})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{\{1\}\{2\}},$ (305)
$I(X_{1};X_{3})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{\{1\}\{3\}},$ (306)
$I(X_{2};X_{3})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{\{2\}\{3\}}$ (307)
To satisfy the non-negativity requirement, we need
$\Pi_{\{1\}\{2\}}=I(X_{1};X_{2})-\Pi_{\{1\}\{2\}\{3\}}\geq 0,$ (308)
$\Pi_{\{1\}\{3\}}=I(X_{1};X_{3})-\Pi_{\{1\}\{2\}\{3\}}\geq 0,$ (309)
$\Pi_{\{2\}\{3\}}=I(X_{2};X_{3})-\Pi_{\{1\}\{2\}\{3\}}\geq 0,$ (310)
which is equivalent to
$0\leq\Pi_{\{1\}\{2\}\{3\}}\leq\min\left(I(X_{1};X_{2}),I(X_{1};X_{3}),I(X_{2};X_{3})\right)$ (311)
The last independent equation can be used in the following form
$I_{3}(X_{1};X_{2};X_{3})=\Pi_{\{1\}\{2\}\{3\}}-\Pi_{s},$ (312)
therefore
$\Pi_{s}=\Pi_{g}=\Pi_{\{1\}\{2\}\{3\}}-I_{3}(X_{1};X_{2};X_{3})\geq 0,$ (313)
The obtained set of conditions is indeed self-consistent, as for any random variables it is true that
$\min\left(I(X_{1};X_{2}),I(X_{1};X_{3}),I(X_{2};X_{3})\right)\geq I_{3}(X_{1};X_{2};X_{3})$ (314)
Table 2: Parthood table for the general 3-variable decomposition
$f$$\{1\}\{2\}\{3\}$$\{1\}\{2\}$$\{1\}\{3\}$$\{2\}\{3\}$$\{12\}\{3\}$$\{13\}\{2\}$$\{23\}\{1\}$$\{1\}$$\{2\}$$\{3\}$$\{12\}$$\{13\}$$\{23\}$$\{123\}$
$\Pi_{\{1\}\{2\}\{3\}}[3]$ 1 1 1 1 1 1 1 1 1 1 1 1 1 1
$\Pi_{s}[2]$ 0 0 0 0 1 1 1 1 1 1 1 1 1 1
$\Pi_{\{1\}\{2\}}[2]$ 0 1 0 0 0 1 1 1 1 0 1 1 1 1
$\Pi_{\{1\}\{3\}}[2]$ 0 0 1 0 1 0 1 1 0 1 1 1 1 1
$\Pi_{\{2\}\{3\}}[2]$ 0 0 0 1 1 1 0 0 1 1 1 1 1 1
$\Pi_{\{1\}}[1]$ 0 0 0 0 0 0 0 1 0 0 1 1 0 1
$\Pi_{\{2\}}[1]$ 0 0 0 0 0 0 0 0 1 0 1 0 1 1
$\Pi_{\{3\}}[1]$ 0 0 0 0 0 0 0 0 0 1 0 1 1 1
$\Pi_{g}[1]$ 0 0 0 0 0 0 0 0 0 0 1 1 1 1
Appendix D Partial information decomposition
The PID atoms are only a subset of our decomposition atoms $\Pi$. Yet, for some systems the PID may contain the full set of atoms. Indeed, when the output is exactly the joint distribution of all inputs, it essentially ’covers’ the whole diagram of the system of inputs. The entropy of inputs completely turns into mutual information about the output.
Lemma 13.
The PID with inputs $X_{1},\dots X_{N}$ and their joint distribution chosen as an output $X_{N+1}=(X_{1},\dots,X_{N})$ contains all information atoms $\Pi$ from the diagram for $X_{1},\dots X_{N+1}$.
Proof.
The PID atoms are by definition all the ones contained in the intersection of the form
$(X_{1}\cup X_{2}\cup\dots\cup X_{N})\cap X_{N+1}$ (315)
By conditions of the lemma, we have
$(X_{1}\cup X_{2}\cup\dots\cup X_{N})\cap X_{N+1}=(X_{1}\cup X_{2}\cup\dots\cup X_{N})\cap(X_{1}\cup X_{2}\cup\dots\cup X_{N})=(X_{1}\cup X_{2}\cup\dots\cup X_{N})=$ (316)
$=(X_{1}\cup X_{2}\cup\dots\cup X_{N})\cup(X_{1}\cup X_{2}\cup\dots\cup X_{N})=(X_{1}\cup X_{2}\cup\dots\cup X_{N})\cup X_{N+1}$ (317)
Hence by lemma 7, all atoms from $(X_{1}\cup X_{2}\cup\dots\cup X_{N})\cup X_{N+1}$, i.e. the whole system, are present in the PID. ∎
A stronger statement can be made that the whole structure of the resulting $N+1$ variable diagram is equivalent to the lesser diagram of just the inputs $X_{1},\dots X_{N}$ with a single extra covering added to each atom to account for the output $X_{N+1}$ ’covering’ the whole system.
Theorem 4.
A solution for the $N+1$ variable decomposition $X_{1},\dots X_{N+1}$ with
$X_{N+1}=X_{1}\cup X_{2}\cup\dots\cup X_{N}$ (318)
defined by a parthood table $f_{\alpha i}$ and a set of atoms $\{\Pi_{i}[c_{i}]\}_{i\in I}$ can be obtained from the decomposition $f^{\prime}_{\alpha^{\prime}j},\{\Pi^{\prime}_{j}[c_{j}^{\prime}]\}_{j\in J}$ of the $N$-variable system $X_{1},\dots X_{N}$ by a transformation
$\forall j\in J\text{ }\exists!i\in I^{\prime}\subset I:\begin{cases}\forall\alpha^{\prime}\hookrightarrow f^{\prime}_{\alpha^{\prime}j}=f_{\alpha^{\prime}i}\\ \Pi_{i}=\Pi^{\prime}_{j}\\ c_{i}=c^{\prime}_{j}+1\end{cases}$ (319)
where the full set of antichains $\{\alpha^{\prime}\}$ in the $N$-variable lattice is a subset of antichains $\{\alpha\}$ in the lattice for $N+1$ variables. The rest of atoms is put to zero
$\forall i\in I\backslash I^{\prime}\hookrightarrow\Pi_{i}=0$ (320)
Proof.
Step 1:
For any antichain $\alpha$, which does not contain index $N+1$
$\alpha=\{k_{11}k_{12}\dots\}\{k_{21}\dots\}\dots\{k_{n1}\dots\},$ (321)
consider several other antichains created by adding the index $N+1$ into one of the brackets
$\alpha^{(-1)}=\alpha,$ (322)
$\alpha^{(0)}=\{N+1\}\{k_{11}k_{12}\dots\}\{k_{21}\dots\}\dots\{k_{n1}\dots\},$ (323)
$\alpha^{(1)}=\{(N+1)k_{11}k_{12}\dots\}\{k_{21}\dots\}\dots\{k_{n1}\dots\},$ (324)
$\alpha^{(2)}=\{k_{11}k_{12}\dots\}\{(N+1)k_{21}\dots\}\dots\{k_{n1}\dots\},$ (325)
$\dots$ (326)
$\alpha^{(n)}=\{k_{11}k_{12}\dots\}\{k_{21}\dots\}\dots\{(N+1)k_{n1}\dots\}$ (327)
It is easy to see that this definition covers all the possible antichains in the redundancy lattice for $N+1$ variables and divides them into disjoint groups. Examining the corresponding partition terms, we see that they are equal within each group
$\Xi_{\alpha^{(0)}}=X_{N+1}\cap\Xi_{\alpha^{(2)}}=\Xi_{\alpha^{(-1)}},$ (328)
$\Xi_{\alpha^{(m>0)}}=(X_{N+1}\cup X_{k_{m1}}\cup\dots)\cap\dots=(X_{k_{m1}}\cup\dots)\cap\dots=\Xi_{\alpha^{(-1)}},$ (329)
By lemma 7 this already guarantees the equivalence of parthood table columns for non-zero atoms.
$\forall\alpha^{(m)},i:\Pi_{i}>0\hookrightarrow f_{\alpha^{(m)}i}=f_{\alpha^{(-1)}i}$ (330)
Let $I^{\prime}$ be the set of indices of all non-zero atoms.
Step 2:
Now, taking the partition terms of the $N$-variable system $X_{1},\dots X_{N}$, we notice that its lattice of antichains $\alpha^{\prime}$ is equivalent to the sublattice of $\alpha^{(-1)}$ from (322). Assuming we know the solution $f^{\prime}_{\alpha^{(-1)}i},\{\Pi^{\prime}_{j}[c_{j}^{\prime}]\}_{j\in J}$ and, thus, the partition terms $\Xi^{\prime}_{\alpha^{(-1)}}$, we substitute
$\forall\alpha^{(-1)},\forall m:-1\leq m\leq|\alpha^{(-1)}|$ (331)
$\Xi^{\prime}_{\alpha^{(-1)}}=\Xi_{\alpha^{(-1)}}=\Xi_{\alpha^{(m)}},$ (332)
$\Xi_{\alpha^{(m)}}=\sum_{j\in J}f^{\prime}_{\alpha^{(-1)}j}\Pi^{\prime}_{j}$ (333)
Let us construct a solution for $N+1$ variables by relating the indices of non-zero atoms of decompositions $\{\Pi^{\prime}_{j}\}_{j\in J}$ and $\{\Pi_{i}\}_{i\in I}$ such that
$\forall\alpha^{(-1)}\hookrightarrow f_{\alpha^{(-1)}i}=f^{\prime}_{\alpha^{(-1)}j},i\in I^{\prime},j\in J$ (334)
Together with (330) this relation fully defines the parthood table $f$ columns, i.e. the set of atoms $\{\Pi_{i}\}$. It is easy to see that this is a bijection between $j\in J$ and $i\in I^{\prime}$. It correctly defines a decomposition for system $X_{1},\dots X_{N+1}$
$\Xi_{\alpha^{(m)}}=\sum_{i\in I}f_{\alpha^{(m)}i}\Pi_{i},$ (335)
[…]
$\forall i\in I\backslash I^{\prime}\hookrightarrow\Pi_{i}=0$ (337)
Step 3:
Most importantly, we need to ensure the validity of the new solution by checking its compliance with all decomposition axioms (198). Here we prove that all newly introduced non-zero atoms indeed exist in the $(N+1)$-variable system, by checking monotonicity for $i\in I^{\prime}$. Notice that
$\alpha^{(l)}\preceq\beta^{(m)}\Rightarrow\alpha^{(-1)}\preceq\beta^{(-1)}$ (338)
Then, (330) combined with the monotonicity condition in the $N$-variable system
$\alpha^{(-1)}\preceq\beta^{(-1)}\Rightarrow\forall i\in I^{\prime}\hookrightarrow f^{\prime}_{\alpha^{(-1)}i}\leq f^{\prime}_{\beta^{(-1)}i}$ (339)
immediately yields
$\alpha^{(l)}\preceq\beta^{(m)}\Rightarrow\forall i\in I^{\prime}\hookrightarrow f_{\alpha^{(l)}i}=f^{\prime}_{\alpha^{(-1)}i}\leq f^{\prime}_{\beta^{(-1)}i}=f_{\beta^{(m)}i}$ (340)
We do not care about the existence of empty atoms $i\in I\backslash I^{\prime}$.
The covering numbers are also adjusted. Notice that
$|\alpha^{(-1)}|=|\alpha^{(m>0)}|=|\alpha^{(0)}|-1$ (341)
Using (198), we find that all coverings $c_{i\in I^{\prime}}$ must be increased by one compared to their pre-images $c^{\prime}_{j}$. Indeed, for the latter we had
$\exists\alpha^{(-1)}:f^{\prime}_{\alpha^{(-1)}j}=1,c^{\prime}_{j}=|\alpha^{(-1)}|,$ (342)
$\forall\beta^{(-1)}:f^{\prime}_{\beta^{(-1)}j}=1\hookrightarrow c^{\prime}_{j}\geq|\beta^{(-1)}|$ (343)
For the resulting covering $c_{i}$ we get
$\exists\alpha^{(0)}:f_{\alpha^{(0)}i}=f_{\alpha^{(-1)}i}=1,c_{i}=|\alpha^{(-1)}|+1=|\alpha^{(0)}|$ (344)
$\forall\beta^{(m)}:f_{\beta^{(m)}i}=f_{\beta^{(-1)}i}=1\hookrightarrow c_{i}\geq|\beta^{(-1)}|+1\geq|\beta^{(m)}|$ (345)
hence
$c_{i}=\max_{\alpha^{(m)}:f_{\alpha^{(m)}i}=1}|\alpha^{(m)}|$ (346)
The information conservation law is also satisfied
$H(X_{N+1})=H(X_{1},\dots X_{N}),$ (347)
$H(X_{1},\dots X_{N})=\sum_{j\in J}\Pi^{\prime}_{j}[c^{\prime}_{j}]$ (348)
$\sum_{k=1}^{N}H(X_{k})=\sum_{j\in J}c^{\prime}_{j}\Pi^{\prime}_{j}[c^{\prime}_{j}]\Leftrightarrow$ (349)
$\Leftrightarrow\sum_{k=1}^{N+1}H(X_{k})=\sum_{j\in J}c^{\prime}_{i}\Pi^{\prime}_{i}[c^{\prime}_{i}]+H(X_{N+1})=\sum_{j\in J}(c^{\prime}_{j}+1)\Pi^{\prime}_{j}[c^{\prime}_{j}]=\sum_{i\in I}c_{i}\Pi_{i}[c_{i}]$ (350)
While we have formulated the overall structure of decomposition, the issue of ambiguity prohibits us from calculating the actual atom sizes. The system [289] is underdetermined: it has 8 equations and 9 variables. This translates to the PID in the classic form of the system from [6] with 3 equations and 4 unknowns
$I(X_{3};X_{1})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{\{1\}\{3\}},$ (351)
$I(X_{3};X_{2})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{\{2\}\{3\}},$ (352)
$I(X_{3};X_{1},X_{2})=\Pi_{\{1\}\{2\}\{3\}}+\Pi_{s}+\Pi_{\{1\}\{3\}}+\Pi_{\{2\}\{3\}}$ (353)
A common method of fixing the remaining degree of freedom is postulating the size of set-theoretic atom $\Pi_{\{1\}\{2\}\{3\}}$ (redundancy) using separate arguments. For the $N$-variable case it is generalized to postulating a new function - redundancy measure - whose values define all set-theoretic atoms and subsequently, all the other atoms of PID. It is defined for a given set of inputs and a output as
$I_{\cap}(X_{N+1}:X_{1};X_{2};\dots X_{N})=\Pi_{\{1\}\{2\}\dots\{N+1\}}$ (354)
In our case this translates to
$I_{\cap}(X_{N+1}:X_{1};X_{2};\dots X_{N})=\hat{H}(X_{1}\cap X_{2}\cap\dots\cap X_{N+1})$ (355)
The redundancy measure can be also used outside of the particular PID it came from. It can be shown that for $N$-variable PID of $X_{1},\dots X_{N}$, choosing a subset of inputs $X_{i_{1}},\dots X_{i_{n}}$ gives the following equality for set-theoretic atoms $\Pi_{\alpha}$
$I_{\cap}(X_{N+1}:X_{i_{1}};X_{i_{2}};\dots X_{i_{n}})=\sum_{\alpha\preceq\{i_{1}\}\{i_{2}\}\dots\{i_{n}\}}\Pi_{\alpha}[|\alpha|]$ (356)
This reformulates the problem of PID ambiguity as a problem of defining a valid redundancy measure[11]. Several axioms were proposed over time to make the results physically sensible [6, 9]:
1) Non-negativity
All information atoms have to be non-negative, since the meaning of negative information is unclear from the physical prospective
$I_{\cap}(X_{N+1}:X_{1};X_{2};\dots X_{N})\geq 0$ (357)
2) Symmetricity
Permutations of the inputs’ indices should not affect the PID
$I_{\cap}(X_{N+1}:\dots X_{i};X_{j};\dots)=I_{\cap}(X_{N+1};\dots X_{j};X_{i};\dots)$ (358)
3) Self-redundancy
If there is a single atom in the PID, it accounts for all the relevant information
$I_{\cap}(X_{N+1}:X_{1})=I(X_{N+1};X_{1})$ (359)
4) Monotonicity
From set-theoretic arguments adding another input should not increase the intersections
$I_{\cap}(X_{N+1}:X_{1};X_{2};\dots X_{N};\tilde{X})\leq I_{\cap}(X_{N+1}:X_{1};X_{2};\dots X_{N})$ (360)
5) Identity
Another condition comes from considering a particular PID with the output being exactly the joint distribution of inputs
$I_{\cap}(X_{1},X_{2}:X_{1};X_{2})=I(X_{1};X_{2})$ (361)
Lemma 14.
Axioms of the redundancy measure
Redundancy measure of the form [354] satisfies all axioms [357-361]. Moreover, it also satisfies strong symmetricity axiom:
$I_{\cap}(X_{N+1}:X_{1};\dots X_{N})=I_{\cap}(X_{i}:X_{1};\dots X_{i-1};X_{N+1};X_{i+1};\dots X_{N}),$ (362)
Proof.
1) Non-negativity follows from the analogous axiom of the measure $\hat{H}$.
2) The intersection $\cap$ operation is associative and commutative.
3)
$I_{\cap}(X_{N+1}:X_{1})=\hat{H}(X_{N+1}\cap X_{1})=I(X_{N+1};X_{1})$ (363)
4)
$X_{N+1}\cap X_{1}\cap\dots\cap X_{N}\supset X_{N+1}\cap X_{1}\cap\dots\cap X_{N}\cap\tilde{X}\Rightarrow$ (364)
$\Rightarrow\hat{H}(X_{N+1}\cap X_{1}\dots\cap X_{N})>\hat{H}(X_{N+1}\cap X_{1}\cap\dots\cap X_{N}\cap\tilde{X})$ (365)
5)
$I_{\cap}(X_{1},X_{2}:X_{1};X_{2})=\hat{H}((X_{1},X_{2})\cap X_{1}\cap X_{2}),$ (366)
$(X_{1},X_{2})=X_{1}\cup X_{2}\supset X_{1}\Rightarrow(X_{1},X_{2})\cap X_{1}=X_{1},$ (367)
$(((X_{1},X_{2})\cap X_{1})\cap X_{2})=X_{1}\cap X_{2},$ (368)
$\hat{H}(X_{1}\cap X_{2})=I(X_{1};X_{2})$ (369)
Within our approach it is possible to bound the redundancy measure from below. We begin by defining a relation $\sim$ between two inputs of a particular PID with output $Y$
$X\sim X^{\prime}\Rightarrow X\cap Y=X^{\prime}\cap Y=(X\cup X^{\prime})\cap Y$ (370)
It is nothing else, but postulating that the only non-zero atoms in the PID $X,X^{\prime},Y$ is redundant information [Fig. 8]. This is not an equivalence relation due to the lack of transitivity.
Figure 8: Diagram of the information atoms for the PID with inputs $X\sim X^{\prime}$.
Lemma 15.
Two inputs contain equivalent information $X\sim X^{\prime}$ about an output $Y$ iff there exist Markov chains $Y\rightarrow X\rightarrow X^{\prime}$ and $Y\rightarrow X^{\prime}\rightarrow X$
Proof.
The existence of said Markov chains is equivalent to
$I((X,X^{\prime});Y)=I(X;Y)=I(X^{\prime};Y)$ (371)
From the subdistributivity axiom we derive that
$X\cap Y\subset(X\cup X^{\prime})\cap Y$ (372)
Taking the difference, we notice that by axioms of the measure it has size zero and therefore is empty
$\left((X\cup X^{\prime})\cap Y\right)\backslash\left(X\cap Y\right)=\emptyset$ (373)
Then by definition
$X\cap Y=(X\cap Y)\cup\emptyset=(X\cup X^{\prime})\cap Y$ (374)
The same is true for $X^{\prime}\cap Y$. All steps are reversible to give the second part of ’iff’ statement. ∎
Theorem 5.
Lower redundancy bound
In the PID with two inputs $X_{1},X_{2}$ and output $X_{3}$, the redundant atom is always larger or equal to the maximum of interaction information over all equivalent inputs:
$\max_{\begin{subarray}{c}X_{1}^{\prime}\sim X_{1}\\ X_{2}^{\prime}\sim X_{2}\end{subarray}}I_{3}(X_{1}^{\prime};X_{2}^{\prime};X_{3})\leq\Pi_{\{1\}\{2\}\{3\}}$ (375)
Proof.
By the definition of input equivalence, we have
$X_{3}\cap X_{1}\cap X_{2}=X_{3}\cap X_{1}^{\prime}\cap X_{2}=X_{3}\cap X_{1}^{\prime}\cap X_{2}^{\prime}$ (376)
Then for any of the systems $X_{1}^{\prime},X_{2}^{\prime}$, the co-information is equal to
$I_{3}(Y;X_{1}^{\prime};X_{2}^{\prime})=\hat{H}(X_{3}\cap X_{1}^{\prime}\cap X_{2}^{\prime})-\Delta\hat{H}$ (377)
The distributivity-breaking term is non-negative, hence
$I_{3}(Y;X_{1}^{\prime};X_{2}^{\prime})\leq\hat{H}(X_{3}\cap X_{1}^{\prime}\cap X_{2}^{\prime})=\hat{H}(X_{3}\cap X_{1}\cap X_{2})=\Pi_{\{1\}\{2\}\{3\}},$ (378)
giving us the statement of the theorem.
We did not use the last condition on the lack of synergy between the equivalent inputs. It is indeed true that the maximum can be taken over a wider collection of systems, however, we would not have any convenient way of finding them similar to lemma 15. ∎